### v1

# Distributional Unlearning: Forgetting Distributions, Not Just Samples

Abstract
Machine unlearning seeks to remove unwanted information from trained models, initially at the individual‐sample level, but increasingly at the level of entire sub‐populations. In many deployments, models must delete whole topical domains to satisfy privacy, legal, or quality requirements, e.g., removing several users’ posts under GDPR or copyrighted web content. Existing unlearning tools remain largely sample-oriented, and straightforward point deletion often leaves enough residual signal for downstream learners to recover the unwanted domain. We introduce distributional unlearning, a data-centric, model-agnostic framework that asks: Given examples from an unwanted distribution and a retained distribution, what is the smallest set of points whose removal makes the edited dataset far from the unwanted domain yet close to the retained one? Using Kullback–Leibler divergence to quantify removal and preservation, we derive the exact Pareto frontier in the Gaussian case and prove that any model retrained on the edited data incurs log-loss shifts bounded by the divergence thresholds. We propose a simple distance-based selection rule satisfying these constraints with a quadratic reduction in deletion budget compared to random removal. Experiments on synthetic Gaussians, Jigsaw Toxic Comments, SMS spam, and CIFAR-10 show 15–72% fewer deletions than random, with negligible impact on retained performance.
1 Introduction
Long-lived models meet short-lived data. Machine learning models often live in production far longer than the data on which they were trained. Over time, significant portions of the training data may become legally or ethically objectionable—think of a large text classifier that scrapes web forums, only to face a GDPR [1] takedown demand requiring the removal of several users’ posting history, or an image‐recognition system that must excise every photo from a now-copyrighted online archive [12]. In these scenarios, deleting a handful of flagged samples is often insufficient: the remaining dataset may still carry the statistical signature of the unwanted domain, allowing downstream models to relearn or exploit it. For example, recent works show that large language models can verbatim reproduce sequences that were removed from their training sets, purely from overlapping context [20]. While “approximate retraining” methods, like influence‐function updates [13] or certified unlearning via convex optimization [14, 25], reduce computational overhead, they do not minimize the number of samples that must be removed to erase a domain’s influence, and hence do not address the core statistical question of sample‐efficiency in unlearning.
Existing research on sample‐level unlearning has mainly focused on per‐record deletions with guarantees on parameter‐space closeness to a retrained model [14, 8]. These approaches typically assume that all unwanted samples have been perfectly identified and focus on how quickly the model parameters can be updated. By contrast, our work asks “which samples—and how many—must be deleted to sever the unwanted distribution’s statistical footprint entirely?” Recent advances in class‑level unlearning [27, 17] and spurious‑feature pruning [21] demonstrate effective forms of domain‑level removal, yet neither provides an explicit, divergence‑based deletion budget for minimal sample removal. Additionally, concept erasure methods, e.g., INLP [24], LEACE [5], provide complementary, model‑internal tools for suppressing unwanted signals post hoc, but because they do not edit the raw data, any downstream retraining can potentially reintroduce those concepts. Finally, robust optimization approaches such as distributionally‐robust risk minimization [22] seek to protect against worst‐case shifts, rather than remove a known shift with provable guarantees on how many and which samples must be deleted.
While these lines of work tackle important aspects of unlearning or robustness, they do not provide a unified, distribution‑level forgetting criterion coupled with minimal deletion guarantees. This gap raises a precise statistical question:
Given samples from an unwanted distribution $p_{1}$ and a retained distribution $p_{2}$, how can we select and minimize the set of points to delete so that the edited empirical distribution is information‐theoretically far from $p_{1}$ but close to $p_{2}$?
To answer this, we introduce distributional unlearning, a data‐centric, model‐agnostic framework that reframes forgetting as a purely statistical operation at the distribution level. We quantify “far” and “close” via the forward Kullback–Leibler (KL) divergence. This choice enables control over downstream log-loss performance with respect to both distributions.
Contributions. We summarize our contributions in tackling the above question as follows:
  • Formal framework: We formalize distributional unlearning via forward-KL constraints; $\alpha$ for removal, $\varepsilon$ for preservation, enabling strong guarantees on downstream log-loss.
  • Pareto frontier & downstream guarantees: We derive the closed-form Pareto frontier of achievable $(\alpha,\varepsilon)$ pairs for Gaussians and characterize its geometry in terms of divergence between the reference distributions for exponential families. We prove that any predictor retrained on the edited dataset incurs log-loss shifts bounded precisely by $\alpha$ and $\varepsilon$.
  • Algorithm & sample efficiency: We study the sample complexity of distributional unlearning with both random and selective removal (of the most divergent samples) mechanisms. We show that selective removal improves sample-efficiency quadratically, in terms of number of removed samples, compared to random removal.
  • Empirical validation: Our selective removal strategy significantly reduces the removal budget versus random deletion. We report on experiments on synthetic Gaussians, Jigsaw toxic comments, SMS spam, and CIFAR-10 images, where we demonstrate 15–72% reductions in deletion budgets over random removal, with little loss in utility on retained domains.
1.1 Related Work
Sample‐level unlearning. Sample‑level unlearning has made impressive strides in fast model updates and formal deletion guarantees [23, 28, 30, 9, 19, 2]. Izzo et al. [15] use influence‐function updates to approximate the effect of deleting a single point without full retraining; Guo et al. [14] provide certified bounds on how close the post‐deletion model is to a scratch‐retrained one; Bourtoule et al. [8] employ data sharding to efficiently erase small batches. However, this class of methods does not address which or how many samples must be removed to eliminate a domain’s overall statistical footprint. Our work complements these techniques by asking not just how to update a model once samples are flagged, but which samples to flag in the first place, and in what quantity.
Model-internal unlearning. Concept erasure methods tackle unwanted attributes in learned features. INLP [24] repeatedly projects out directions predictive of a protected attribute; counterfactual augmentation [16] synthesizes targeted data to sever causal links; adversarial training [11] trains encoders to remove specific signals. These operate post‐hoc on a fixed model’s representations—ideal for fairness use-cases such as removing gender or sentiment—but they rely on white-box access and tailor to one model at a time. Where concept erasure edits model representations, we edit the data, guaranteeing forgetting for downstream models.
Domain adaptation, robustness, and coresets. Domain adaptation theory bounds target error by source error plus a divergence between domains [6]. Our work flips and complements this paradigm: we intentionally increase divergence from an unwanted source $p_{1}$ while controlling proximity to a desired reference $p_{2}$. Distributionally‐robust optimization [22] protects against all shifts within a divergence ball, whereas we target the removal of one specific shift. Coreset and importance‐sampling methods [26] select representative subsets to approximate a distribution; we invert that idea to remove the most representative samples of the unwanted component, while preserving another.
2 Distributional Unlearning: Definition and Implications
We consider probability distributions over an input space $\mathcal{X}$, and denote by $\mathcal{P}$ a class of distributions on $\mathcal{X}$. We focus on two distributions:
  • $p_{1}\in\mathcal{P}$: the target distribution to forget,
  • $p_{2}\in\mathcal{P}$: the reference distribution to preserve.
The goal is to construct a new distribution $p\in\mathcal{P}$ that removes the statistical influence of $p_{1}$ while retaining the properties of $p_{2}$. We formalize this via Kullback–Leibler (KL) divergence, which we recall for two absolutely-continuous distributions $q,p$ on $\mathcal{X}$ is $\mathrm{KL}(q\|p){\coloneqq}\int_{\mathcal{X}}q(x)\,\log\!\frac{q(x)}{p(x)}\,dx.$ Throughout this section, we defer all proofs to Appendix A.
Definition 1 ($(\alpha,\varepsilon)$-Distributional Unlearning).
For tolerances $\alpha,\varepsilon>0$, a distribution $p\in\mathcal{P}$ satisfies $(\alpha,\varepsilon)$-distributional unlearning with respect to $(p_{1},p_{2})$ if:
[…]
The first inequality forces the edited data to be information‑theoretically distant from the population we wish to forget. The second inequality upper bounds collateral damage to the population we preserve. We adopt the forward KL divergence as it directly controls the expected log-loss under the edited distribution $p$, which is critical for bounding downstream predictive performance (Prop. 2). [^1] Reverse KL or Wasserstein could be used to enforce other tail behaviors, which we leave for future work.
2.1 Feasibility and the Pareto Frontier
The pair $(\alpha,\varepsilon)$ captures a trade‑off: how far we can move from $p_{1}$ while remaining close to $p_{2}$. To understand which $(\alpha,\varepsilon)$ pairs are jointly achievable, we characterize the feasible region and its boundary. Formally, the Pareto frontier $\mathrm{PF}(p_{1},p_{2};\mathcal{P})$ consists of those pairs $(\alpha,\varepsilon)$ for which no strictly better trade-off exists: there is no $p^{\prime}\in\mathcal{P}$ satisfying $\mathrm{KL}(p_{1}\|p^{\prime})\geq\alpha^{\prime}$ and $\mathrm{KL}(p_{2}\|p^{\prime})\leq\varepsilon^{\prime}$ with $\alpha^{\prime}>\alpha$ and $\varepsilon^{\prime}<\varepsilon$. That is, every point on the frontier is optimal in the sense that one objective cannot be improved without worsening the other.
The following result gives the closed-form Pareto frontier in the Gaussian case:
Proposition 1 (Pareto Frontier).
Let $p_{1},p_{2}$ be two distributions in $\mathcal{P}$, the class of Gaussian distributions with shared covariance, with $\mathrm{KL}(p_{1}\|p_{2})<\infty$. The Pareto frontier of $(\alpha,\varepsilon)$ values achievable by any $p\in\mathcal{P}$ is given by:
$\mathrm{PF}(p_{1},p_{2};\mathcal{P})=\left\{\left(\alpha,\left(\sqrt{\alpha}-\sqrt{\mathrm{KL}(p_{1}\|p_{2})}\right)^{2}\right):\alpha\geq\mathrm{KL}(p_{1}\|p_{2})\right\}.$
This frontier characterizes the minimal preservation loss $\varepsilon$ required to achieve a given removal level $\alpha$. In particular, the value $\mathrm{KL}(p_{1}\|p_{2})$ plays a critical role: no distribution can forget more than this amount ($\alpha{>}\mathrm{KL}(p_{1}\|p_{2})$) while remaining arbitrarily close to $p_{2}$. The shape of the frontier reflects how intertwined $p_{1}$ and $p_{2}$ are. While this curve was derived analytically for Gaussians, the same removal–preservation tradeoff arises more broadly, e.g., regular exponential families (Proposition 3). Empirically, the frontier matches closely in synthetic Gaussian experiments (Fig. 1), confirming both its shape and the threshold behavior.
2.2 Implications for Downstream Prediction
We now connect distributional unlearning to predictive performance in supervised learning. Consider a predictor $h:\mathcal{X}\to\Delta(\mathcal{Y})$, where $\mathcal{X}$ is the input space and $\Delta(\mathcal{Y})$ the probability simplex over label space $\mathcal{Y}$, trained on a distribution $p$ over $\mathcal{X}\times\mathcal{Y}$ that satisfies $(\alpha,\varepsilon)$-unlearning with respect to $(p_{1},p_{2})$. We study how $h$ performs under the true data-generating distributions $p_{1}$ and $p_{2}$.
Let $\ell(y,q)\coloneqq-\log q(y)$, for $y\in\mathcal{Y},q\in\Delta(\mathcal{Y})$, denote the log-loss. Define the expected loss under $p$ as $\mathcal{L}(h;p)\coloneqq\mathbb{E}_{(x,y)\sim p}[\ell(y,h(x))]$. Then, for any class of distributions $\mathcal{P}$, we have:
Proposition 2.
Let $h$ minimize $\mathcal{L}(h;p)$, and let $h_{1}$, $h_{2}$ be optimal predictors under $p_{1},p_{2}\in\mathcal{P}$, respectively. If $p$ satisfies $(\alpha,\varepsilon)$-distributional unlearning with respect to $(p_{1},p_{2})$, then:
$\mathcal{L}(h;p_{1})-\mathcal{L}(h_{1};p_{1})\geq\alpha-\delta_{1},$ $\mathcal{L}(h;p_{2})-\mathcal{L}(h_{2};p_{2})\leq\varepsilon-\delta_{2},$ (2)
where $\delta_{1}:=\mathrm{KL}(p_{1,X}\|p_{X}),\delta_{2}:=\mathrm{KL}(p_{2,X}\|p_{X})$ denote the marginal KL divergence over inputs.
These bounds show that distributional unlearning guarantees increased loss under the forgotten distribution and bounded degradation under the preserved one. In this sense, the $(\alpha,\varepsilon)$-distributional unlearning framework provides meaningful control over downstream predictive behavior. Regarding the extra marginal KL term in the first inequality, which quantifies divergence on input distributions, the data‑processing inequality gives $\mathrm{KL}\!\bigl{(}p_{1,X}\|p_{X}\bigr{)}\!\leq\!\mathrm{KL}(p_{1}\|p)$, hence this extra term is always bounded by the same $\alpha$ we already control. A similar term appears in the second inequality, but can only improve the preservation bound.
3 Algorithms and Sample Complexity
We now instantiate the distributional unlearning framework in a concrete setting where we only have access to samples from $p_{1}$ and $p_{2}$, and we must decide which points to remove and how many in order to achieve $(\alpha,\varepsilon)$‑distributional unlearning. In this section, we (i) introduce two deletion strategies—random removal and selective removal—and (ii) derive high‑probability bounds on the number of samples that must be deleted under each.
In the following, we focus on the univariate Gaussian case, since it captures the essential phenomena. That is, denoting $\mathcal{P}\coloneqq\left\{\mathcal{N}{(\mu,\sigma^{2})}\colon\mu\in\mathbb{R}\right\}$, where $\sigma>0$, we have $p_{1},p_{2}\in\mathcal{P}$. We consider $n_{1}$ i.i.d. samples from $p_{1}$ and $n_{2}$ from $p_{2}$. We let $0\leq f\leq n_{1}$ denote the number of points removed from the $p_{1}$ samples. We defer all proofs to Appendix A.
Mechanism Sample Complexity
Removal Preservation
Random Removal (Thm. 1) $n_{1}\left(1-\sqrt{1-\alpha}\right)$ $n_{1}\left(1-{\sqrt{\varepsilon}}\right)$
Selective Removal (Thm. 2) $n_{1}\left(1-\left(1-\alpha\right)^{1/4}\right)$ $n_{1}\left(1-\varepsilon^{1/4}\right)$
Table 1: Summary of simplified sample complexity bounds: the formulas give the minimal number of $p_{1}$ samples (out of $n_{1}{\geq}1$) to remove in order to achieve $(\alpha,\varepsilon)$‐distributional unlearning, and show the quadratic improvement of selective removal over random in $\varepsilon,\alpha$. We focus on dependences on $\alpha,\varepsilon\in(0,1)$ and ignore constants, $\mathrm{KL}(p_{1}\|p_{2})$, and $\tfrac{n_{2}}{n_{1}}$, and assume that $n_{2}$ is large enough. Full expressions and assumptions are given in Corollary 8 in the appendix.
3.1 Random Removal
We begin with a baseline deletion strategy that treats every sample equally, deleting $f$ points chosen uniformly at random from the $n_{1}$ samples of $p_{1}$. For clarity, we first state the random removal procedure formally below:
Algorithm (Random Removal).
  1. 1.
    Randomly select $f$ out of the $n_{1}$ samples of $p_{1}$ without replacement.
  2. 2.
    Remove those $f$ samples.
  3. 3.
    Re‑fit $\mathcal{N}(\hat{\mu},\sigma^{2})$ by MLE on the remaining data.
The following theorem provides a finite-sample guarantee for achieving $(\alpha,\varepsilon)$-distributional unlearning using random removal with a deletion budget $f$.
Theorem 1 (Random Removal).
Let $p_{1},p_{2}\in\mathcal{P}$ and $\delta\in(0,1)$. We observe $n_{1}$ samples from $p_{1}$ and $n_{2}$ samples from $p_{2}$, and randomly remove $f$ samples from $p_{1}$ before fitting. Then with probability at least $1-\delta$, the resulting MLE distribution satisfies $(\alpha,\varepsilon)$-distributional unlearning with:
$\alpha$ $\geq\left(\frac{1}{2}-3\left(\frac{n_{1}-f}{n_{2}}\right)^{2}\right)\mathrm{KL}(p_{1}\parallel p_{2})-\frac{3\ln(4/\delta)}{2n_{2}}\left(1+\frac{n_{1}-f}{n_{2}}\right),$
$\varepsilon$ $\leq 3\left(\frac{n_{1}-f}{n_{2}}\right)^{2}\mathrm{KL}(p_{1}\parallel p_{2})+\frac{3\ln(4/\delta)}{n_{2}}\left(1+\frac{n_{1}-f}{n_{2}}\right).$
This result shows that random removal can achieve both forgetting and preservation as long as $f$ is large enough, so that the term $\tfrac{n_{1}-f}{n_{2}}$, i.e., ratio of number of non-deleted $p_{1}$ samples to $p_{2}$ samples, is small enough. First, we remark that the last term in both inequalities can be ignored if $n_{2}$ is large enough compared to $\log(1/\delta)$. Next, the removal guarantee improves linearly in $\mathrm{KL}(p_{1}\|p_{2})$, but quadratically in $\tfrac{n_{1}-f}{n_{2}}$, indicating diminishing returns as more data is deleted. Note that preservation improves as $n_{2}$ increases, since the estimation of $p_{2}$ becomes more accurate, reducing collateral damage. While conceptually simple, this method does not prioritize which samples contribute most to the divergence from $p_{2}$, and may thus be inefficient in certain regimes.
3.2 Selective Removal
We next consider a selective removal strategy that uses knowledge of the data to delete the most distinguishable samples from $p_{2}$, specifically the lowest likelihood under the empirical estimation of the preserved distribution $p_{2}$. We outline the exact procedure below:
Algorithm (Selective Removal).
  1. 1.
    Compute the mean $\hat{\mu}_{2}$ of the $n_{2}$ samples from $p_{2}$.
  2. 2.
    For each of the $n_{1}$ samples $x_{i}$ from $p_{1}$, compute the score $s_{i}=|x_{i}-\hat{\mu}_{2}|$.
  3. 3.
    Delete the $f$ samples with the largest scores $s_{i}$.
  4. 4.
    Re‑fit $\mathcal{N}(\hat{\mu},\sigma^{2})$ by MLE on the remaining data.
The following theorem provides a finite-sample guarantee for achieving $(\alpha,\varepsilon)$-distributional unlearning using selective removal with a deletion budget $f$.
Theorem 2 (Selective Removal).
Let $p_{1},p_{2}\in\mathcal{P}$ and $\delta\in(0,1)$. Let $f$ samples from $p_{1}$ be removed according to Selective Removal. Then with probability at least $1-\delta$, the resulting estimate satisfies $(\alpha,\varepsilon)$-distributional unlearning with:
$\alpha$ $\geq\frac{1}{2}\mathrm{KL}(p_{1}\parallel p_{2})-\frac{1}{2}\left(\frac{n_{1}-f}{n_{2}}\right)^{2}g^{-1}\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}};\mathrm{KL}(p_{1}\parallel p_{2})\Bigr{)}^{2}-{\frac{\ln(4/\delta)}{n_{2}}},$
$\varepsilon$ $\leq\left(\frac{n_{1}-f}{n_{2}}\right)^{2}g^{-1}\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}};\mathrm{KL}(p_{1}\parallel p_{2})\Bigr{)}^{2}+\frac{2\ln(4/\delta)}{n_{2}},$
where $g(u;\kappa)\coloneqq\Phi(u-\sqrt{2\kappa})+\Phi(u+\sqrt{2\kappa})-1$, for $u,\kappa>0$, and $\Phi$ is the standard normal CDF.
While this result also shows that larger deletion budgets improve both removal and preservation, the main difference with random removal is that the quadratic factor $(\tfrac{n_{1}-f}{n_{2}})^{2}$ is amplified by a multiplicative factor involving the inverse of $g(\cdot~{};\kappa)$, which represents the CDF of the folded normal shifted by divergence $\kappa=\mathrm{KL}(p_{1}\|p_{2})$. Therefore, for $p\in(0,1)$, $g^{-1}(p;\kappa)$ is the $p$-th quantile of the folded normal shifted by the aforementioned divergence $\kappa$. Intuitively, this term arises as scores computed by selective removal, which are of the form $s_{i}=|x_{i}-\hat{\mu}_{2}|$ with $x_{i}$ being a sample from $p_{1}$, are distributed following a folded normal. Thus, we are effectively selecting samples from $p_{1}$ whose distance to the mean of $p_{2}$ is at most the $\approx(1-\frac{f}{n_{1}})$-th quantile of the aforementioned folded normal, which is fortunately a decreasing function in $f$. That is, this term strictly amplifies the quadratic decrease in $f$, which was the best we could previously obtain with random removal.
Simplified bounds. While the above theorem is general and gives exact quantities, the term involving $g^{-1}(\cdot~{};\kappa)$ can be upper bounded with a simpler analytical expression. As we mentioned previously, this term represents a quantile of the folded normal distribution. We can bound this quantile with a linear approximation, when the divergence $\mathrm{KL}(p_{1}\|p_{2})$ is small enough, i.e., $p_{1}$ and $p_{2}$ are similar, and $\tfrac{f}{n_{1}}$ is large enough, i.e., we delete enough samples. The result of this simplification is given in Table 1, and further developed in Corollary 8. Essentially, focusing only on the dependence on $\alpha\in(0,1)$ for removal, selective removal only requires $\mathcal{O}(n_{1}(1-\sqrt{1-\alpha}))$, while random removal requires $\mathcal{O}(n_{1}(1-(1-\alpha)^{1/4}))$. Analogous conclusions hold for $\varepsilon$ for preservation. Therefore, we show that selective removal strictly improves, at least quadratically, over random removal with respect to $\varepsilon,\alpha$ in low-divergence regimes.
We empirically validate the improvement over random removal in Figure 2. Specifically, as discussed earlier, when $\mathrm{KL}(p_{1}\|p_{2})$ is small (top left plot), selective removal significantly improves over random removal. When the distributions are very far apart (top right plot), the difference narrows and both mechanisms perform comparably.
4 Empirical Validation
We now evaluate distributional unlearning empirically across synthetic as well as real-world data. Our goals are threefold:
  1. 1.
    Compare the (deletion) sample-efficiency of random versus selective deletion strategies.
  2. 2.
    Assess how distributional unlearning shapes downstream predictive performance.
  3. 3.
    Validate the theoretical Pareto frontier of achievable $(\alpha,\varepsilon)$ values.
We consider four qualitatively different settings: synthetic Gaussians, Jigsaw toxic-comment moderation, SMS spam filtering, and image classification on CIFAR-10. In each case, we define a distribution $p_{1}$ to forget and a distribution $p_{2}$ to preserve, apply several deletion methods to remove samples from $p_{1}$, and measure the resulting trade-off in statistical proximity and model performance. All results are averaged over multiple random seeds. We defer experimental details to Appendix B.
Domain Target on $p_{1}$ Random Selective Saving
Gaussians (low) $\mathrm{KL}(p_{1}\|p)$ 65 18 72 %
Gaussians (high) $\mathrm{KL}(p_{1}\|p)$ 65 50 23 %
Jigsaw toxic comments Recall 100 85 15 %
SMS Spam Recall 90 75 17 %
[…]
Table 2: Deletion budget (%) needed to reach half of the initial value of the removal metric (no deletion) on each dataset. “Selective”=best‑performing selective removal score; “Saving”=relative reduction versus random deletion. Gaussians (low) and (high) are the scenarios of the top leftmost and rightmost plots of Fig. 2, respectively.
Results overview. We summarize our main findings in Table 2 for convenience, before detailing them in the next sections. Across Gaussians, toxic text, short‑message spam, and natural images, the same selective removal recipe cuts the deletion budget by $15\!-\!72\,\%$ relative to random removal while reducing accuracy or recall on the forget distribution by at least half. The gains are largest in the low‑divergence Gaussian regime ($72\,\%$), where theory predicts the greatest advantage for selective deletion. The same trends persist in high‑diversity real data ($15\!-\!38\,\%$) confirming that distributional unlearning controls downstream predictive performance as predicted by Proposition 2. Crucially, we achieve these savings without harming utility on the retained distribution. Table 2 distills the central empirical message: selective removal delivers the desired distributional unlearning guarantees with substantially less data removal than naïve approaches consistently across domains and models.
4.1 Synthetic Gaussians: frontier verification and sample efficiency
We begin with a synthetic setting to evaluate distributional unlearning under controlled conditions. Let $p_{1}=\mathcal{N}(0,1)$ and $p_{2}=\mathcal{N}(\mu,1)$ for varying $\mu\in\{0.5,2.5,5\}$, which induces increasing divergence $\mathrm{KL}(p_{1}\|p_{2})$. For each configuration, we draw $n_{1}=n_{2}=1000$ samples from $p_{1}$ and $p_{2}$, respectively, and examine how different deletion strategies affect the post-edit distribution. We implement two mechanisms: random removal, which deletes a uniform subset of $p_{1}$ points, and selective removal, which prioritizes those samples with largest deviation from the empirical mean of $p_{2}$. After deleting $f$ samples, we re-fit a Gaussian $\mathcal{N}(\hat{\mu},1)$ on the retained $p_{1}$ and $p_{2}$ samples via maximum likelihood, yielding a post-edit distribution ${p}$. We then compute forward KL divergences $\alpha=\mathrm{KL}(p_{1}\|{p})$ and $\varepsilon=\mathrm{KL}(p_{2}\|{p})$ to quantify forgetting and preservation, respectively.
Figure 1: Synthetic Gaussians. The empirical frontier aligns with the theoretical prediction.
Pareto frontier. Figure 1 confirms that the empirical $(\alpha,\varepsilon)$ trade-off closely matches the theoretical Pareto frontier derived in Proposition 1. To plot feasible empirical trade-offs, we set $p=\mathcal{N}(\mu,1)$ and vary $\mu\in\mathbb{R}$, while $p_{1}=\mathcal{N}(0,1)$ and $p_{2}=\mathcal{N}(2,1)$ so that $\mathrm{KL}(p_{1}\|p_{2})=2$. The latter quantity is the threshold predicted by the theory, and validated by Figure 1. Indeed, feasible trade-offs whose removal divergence $\alpha$ is below this threshold are pareto-suboptimal. They are dominated by the trade-off $(\alpha=\mathrm{KL}(p_{1}\|p_{2}),\varepsilon=0)$, which can be achieved with the choice of distribution $p=p_{2}$.
Figure 2: Synthetic Gaussians. Selective removal consistently requires fewer deletions, especially when $\mathrm{KL}(p_{1}\|p_{2})$ is small (left), for the same removal and preservation target as random removal. In high-divergence regimes (right), the gap between methods shrinks, as predicted by the theory.
Sample efficiency. We next compare the efficiency of the two removal strategies in the finite-sample case analyzed in Section 3. In Figure 2, we plot achieved $\alpha$ as a function of the number of $p_{1}$ samples removed. Selective removal reaches higher forgetting levels with fewer deletions than random removal, especially when $\mathrm{KL}(p_{1}\|p_{2})$ is small ($\mu_{2}=0.5$, top left plot). For example, to reach $0.06$ nats of removal divergence, i.e., half of that obtained by removing all samples, selective removal requires 5×5$\times$5 × less samples than random removal. Analogous trends hold for preservation (bottom left plot): selective removal more effectively preserves the reference distribution $p_{2}$ throughout. The remaining plots show similar trends when increasing the divergence $\mathrm{KL}(p_{1}\|p_{2})$ by increasing $\mu_{2}$. As predicted in theory, selective removal offers the greatest savings when $p_{1}$ and $p_{2}$ are close and diminishes as the distributions diverge.
4.2 Jigsaw Toxic: forgetting a toxic sub‑population
We now consider the Jigsaw toxic comments dataset [10]. Let $p_{1}$ be the set of comments that contain any of five high‑frequency profanity keywords (see Appendix B). This sub‑population accounts for $8.6\,\%$ of the training corpus. All remaining comments form the reference distribution $p_{2}$. For each deletion budget $f$ corresponding to removal fractions $\!\{0,0.05,\dots,1\}$, we remove $f$ comments from $p_{1}$ ranked by four scoring rules computed based on TF-IDF embeddings, plus a random baseline:
  1. 1.
    cos‑mu2: cosine distance to the $p_{2}$ mean only;
  2. 2.
    lr‑cos: cosine margin between the comment and the mean vectors of $p_{2}$ vs. $p_{1}$;
  3. 3.
    knn‑ratio: local $k$‑NN density ratio ($k{=}10$);
  4. 4.
    tfidf‑norm: comment $\ell_{2}$ length; and
  5. 5.
    random: uniformly random baseline.
The first heuristic cos‑mu2 above is a proxy for the selective removal method introduced in Section 3, where we use cosine distance between TF-IDF embeddings instead of Euclidean distance. The second heuristic lr‑cos is a simple extension, and stands for a proxy of a likelihood ratio score, i.e., we remove samples most distinguishable from $p_{2}$ while being representive of $p_{1}$. After deletion we re‑train a logistic–TF-IDF model on the kept data. We report Recall@$p_{1}$ (higher means less forgetting) and macro‑F1@$p_{2}$ (higher means stronger preservation).
(a) Recall on profane comments ($p_{1}$, forgotten).
(b) Macro-F1 on non-profane comments ($p_{2}$, kept).
Figure 3: Jigsaw Toxic Comments. Impact of removing profane comments on Jigsaw Toxic. Left: recall on the to‑be‑forgotten set $p_{1}$; right: macro‑F1 on the retained set $p_{2}$. Utility is almost unchanged up to $60\%$ deletion; marked forgetting appears only around $80\%$ deletion, with lr‑cos showing the steepest drop. Error bars: ±$\pm 1$  standard error over five randomness seeds.
Findings. We report our main findings on our Jigsaw task in Figure 3. Across all heuristics, Recall@ $p_{1}$ remains essentially flat around $0.88$ until large budgets; a pronounced forgetting effect appears only at $f\!\geq 0.8|p_{1}|$, where lr‑cos decreases recall by nearly $20$ percentage points (Fig. 3(a)). Crucially, macro‑F1@ $p_{2}$ is almost unchanged up to $f\!=\!0.6|p_{1}|$ (Fig. 3(b)) and declines gradually afterwards, indicating that removing up to $60\,\%$ of the profane comments has negligible impact on overall utility. The sharp drop at full deletion confirms the theoretical pareto frontier prediction that excessive removal targets must harm performance on the retained distribution. This can be naturally explained on this toxic comment detection task, since $p_{1}$ the distribution of profane comments, carries task‑relevant signal. By contrast, random removal requires removing most of $p_{1}$, the distribution of profane comments, to reach the same removal level of selective methods. Specifically, random removal requires removing $95\%$ of $p_{1}$ to achieve around $0.7$ recall, which means that LR-cos saves around $16\%$ and cos-mu2 saves $11\%$ in terms of deletion budget compared to random removel for the same Recall@ $p_{1}$ target. All other methods have a similar performance to random removal.
4.3 SMS Spam: a content-moderation unlearning task
We revisit the UCI SMS Spam Collection [3], treating the spam class ($p_{1}$) as information to forget and the ham class ($p_{2}$) as information to preserve. Messages are vectorised with TF–IDF features. Deletion budgets again span $5\%$ to $100\%$ of the spam slice in 5-point increments. We compare the same five scoring rules as before (cos-mu2, lr-cos, knn-ratio, tfidf-norm, random) and average results over ten random seeds. Metrics reported are recall on $p_{1}$ and macro-F1 on $p_{2}$.
(a) Recall on spam ($p_{1}$, forgotten).
(b) Macro-F1 on ham ($p_{2}$, kept).
Figure 4: SMS Spam. The likelihood-ratio score (lr-cos) pushes spam recall below $0.6$ after deleting $70\%$ of spam, whereas random deletion needs nearly $90\%$ removal to reach the same point. Ham performance remains almost flat (${<}0.004$ absolute change) until the final 100 % budget, affirming the tight preservation guarantee. Error bars: ±$\pm 1$  standard error over ten seeds.
[…]
4.4 CIFAR‑10: privacy‑motivated forgetting of an entire class
We treat the CIFAR‑10 [18] cat category as a privacy‑sensitive sub‑population. Deleting a few individual images is insufficient; their aggregate statistics would still influence the model. Instead, we rank every cat image with three distance‑based scores (lr-maha, maha-mu2, knn-ratio), and a random baseline. The former two methods are the direct equivalents of lr-cos and cos-mu2, by using the pretrained ResNet18 features to compute scores instead instead of the TF-IDF embeddings, and using the Mahalanobis distance [^2] The Mahalanobis distance of vector $x$ to probability distribution $p$, of mean $\mu$ and covariance $\Sigma$, is: $d(x,p)\coloneqq\sqrt{(x-\mu)^{\top}\Sigma^{-1}(x-\mu)}$. We estimate $\mu$ and $\Sigma$ empirically on the retained distribution $p_{2}$. instead of cosine distance. We delete the top score-ranked samples for each deletion budget, re‑train a CNN for ten epochs, and report accuracy on the cat test set and accuracy on the other nine classes test set.
(a) Accuracy on cat images ($p_{1}$, forgotten).
(b) Accuracy on non-cat images ($p_{2}$, kept).
Figure 5: CIFAR‑10 images. Removing cat images suppresses accuracy on that class (left) while leaving accuracy on the retained nine classes essentially unchanged (right, ${<}0.03$ variation). No substantial removal is observed until $50\%$ deletion, before selective removal strategies lr-maha and maha-mu2 outperform random removal. Error bars: ±$\pm 1$  standard error over thirty seeds.
Findings. Figure 5 shows that the cat footprint persists in terms of accuracy, which is problematic for privacy scenarios and motivates distributional unlearning, until ${\geq}50\%$ data is removed (Fig. 5(a)), yet the model’s utility on the other classes remains stable (Fig. 5(b)) at around $82\%$. Selective removal strategies reach a given accuracy threshold, on the cat class, with substantially fewer deletions than random. Specifically, lr-maha halves the initial accuracy by deleting half of the cat images, which is 1.6×1.6$\times$1.6 × more sample-efficient than random removal. Also, maha-mu2 reaches the same accuracy threshold with 1.3×1.3$\times$1.3 × fewer deletions than random removal. These results underscore the value of distributional unlearning: strong class‑level forgetting is achieved long before every single cat image is deleted, and with minimal collateral loss on the retained distribution.
5 Conclusion and Discussion
Modern machine unlearning demands more than individual record deletion: in many practical settings, models must forget entire subpopulations whose statistical imprint persists even after standard sample-level removal. We introduced distributional unlearning, a framework for erasing the influence of an unwanted data distribution while preserving another, via forward-KL divergence thresholds $(\alpha,\varepsilon)$ that balance removal and preservation. We derived the exact Pareto frontier in the Gaussian case and proved that these constraints induce tight log-loss guarantees for downstream predictors. Analyzing both random deletion and divergence-informed selective removal, we showed that the latter achieves unlearning with substantially fewer deletions, especially when $p_{1}$ and $p_{2}$ are close. Experiments across synthetic and real-world domains validated our theory, showing up to 72% savings in deletion budget with no loss in retained utility.
While our approach operates by editing samples drawn from structured distributions, the core idea extends naturally to representation-level interventions, structured latent-variable models, or causal inference frameworks. Distributional unlearning may also enhance fairness, robustness, or interpretability by targeting and removing harmful subpopulation signals. By operating at the level of data distributions rather than model internals, it complements existing unlearning and debiasing techniques in a model-agnostic fashion. As a general-purpose statistical primitive, it offers a new lens for thinking about modularity and controllability in data-centric learning pipelines.
Appendix A Proofs
A.1 Pareto Frontier
See 1
Proof.
For simplicity, we consider univariate Gaussians with share variance. For $d$‑dimensional Gaussians with covariance $\Sigma\in\mathbb{R}^{d\times d}$, the same result holdsafter replacing squared error by the Mahalanobis distance $\|x-\mu\|_{\Sigma^{-1}}^{2}$ (see Proposition 3).
Let $p=\mathcal{N}(\mu,\sigma^{2})\in\mathcal{P}$. Since all distributions in $\mathcal{P}$ share the same variance, the KL divergence from $p_{i}$ to $p$ is:
$\mathrm{KL}(p_{i}\|p)=\frac{(\mu_{i}-\mu)^{2}}{2\sigma^{2}},\quad i=1,2.$
Fix $\alpha\geq\mathrm{KL}(p_{1}\|p_{2})$. We want to compute the minimal possible $\varepsilon$ achievable under the constraint $\mathrm{KL}(p_{1}\|p)\geq\alpha$. Define this minimum as:
$\varepsilon_{\star}(\alpha)\coloneqq\min_{\mu\in\mathbb{R},~{}(\mu-\mu_{1})^{2}\geq 2\sigma^{2}\alpha}\frac{(\mu_{2}-\mu)^{2}}{2\sigma^{2}}.$
This is a one-dimensional quadratic minimization problem subject to a quadratic inequality constraint. The feasible set is:
$\mu\in(-\infty,\mu_{1}-\sigma\sqrt{2\alpha}]\cup[\mu_{1}+\sigma\sqrt{2\alpha},\infty).$
We minimize $(\mu_{2}-\mu)^{2}$ over this set. This yields two cases:
  • If $\mu_{2}\in[\mu_{1}-\sigma\sqrt{2\alpha},\mu_{1}+\sigma\sqrt{2\alpha}]$, then the closest feasible points are the endpoints. The minimizing value of $\mu$ is:
    $\mu=\mu_{1}+\operatorname{sign}(\mu_{2}-\mu_{1})\cdot\sigma\sqrt{2\alpha},$
    and the resulting divergence is:
    $\varepsilon_{\star}(\alpha)=\frac{(\mu_{2}-\mu_{1}-\operatorname{sign}(\mu_{2}-\mu_{1})\cdot\sigma\sqrt{2\alpha})^{2}}{2\sigma^{2}}.$
  • If $\mu_{2}$ already lies in the feasible set, i.e., $|\mu_{2}-\mu_{1}|\geq\sigma\sqrt{2\alpha}$, then we can choose $\mu=\mu_{2}$, yielding $\varepsilon_{\star}(\alpha)=0$.
Thus, for all $\alpha\geq 0$:
$\varepsilon_{\star}(\alpha)=\frac{\left[\left(|\mu_{2}-\mu_{1}|-\sigma\sqrt{2\alpha}\right)_{+}\right]^{2}}{2\sigma^{2}},$
where $(x)_{+}=\max\{x,0\}$. Let $\Delta\coloneqq|\mu_{2}-\mu_{1}|$, and recall that $\mathrm{KL}(p_{1}\|p_{2})=\frac{\Delta^{2}}{2\sigma^{2}}$. Then:
$\Delta=\sigma\sqrt{2\mathrm{KL}(p_{1}\|p_{2})}.$
Substituting into $\varepsilon_{\star}(\alpha)$:
$\varepsilon_{\star}(\alpha)=\left(\sqrt{\alpha}-\sqrt{\mathrm{KL}(p_{1}\|p_{2})}\right)^{2},\quad\text{for }\alpha\geq\mathrm{KL}(p_{1}\|p_{2}).$
Finally, note that any pair $(\alpha,\varepsilon)$ with $\alpha<\mathrm{KL}(p_{1}\|p_{2})$ satisfies $\varepsilon_{\star}(\alpha)=0$, hence is dominated by $(\mathrm{KL}(p_{1}\|p_{2}),0)$. Therefore, the Pareto optimal points are exactly:
$\left\{\left(\alpha,\left(\sqrt{\alpha}-\sqrt{\mathrm{KL}(p_{1}\|p_{2})}\right)^{2}\right):\alpha\geq\mathrm{KL}(p_{1}\|p_{2})\right\},$
as claimed. ∎
Next, we show that a qualitatively similar result holds more generally for any exponential‑family member.
Proposition 3 (Pareto Frontier–Exponential Families).
Let $(\mathcal{X},\mu)$ be a measurable space and let
$\mathcal{P}=\bigl{\{}\,p_{\theta}(x)=h(x)\exp(\theta^{\top}T(x)-A(\theta))~{}\colon~{}\theta\in\Theta\subset\mathbb{R}^{d}\bigr{\}}$
[…]
  1. (i)
    The Pareto frontier for points in $\mathcal{P}$ is
    $\mathrm{PF}(p_{1},p_{2};\mathcal{P})=\Bigl{\{}\bigl{(}\alpha,\,v(\alpha)\bigr{)}:\alpha\geq\mathrm{KL}(p_{1}\|p_{2})\Bigr{\}},$
    where $v(\alpha)=\mathrm{KL}(p_{2}\|p_{1})+\alpha+\frac{1}{\lambda^{*}-1}(\theta_{2}-\theta_{1})^{\top}(\operatorname{{\mathbb{E}}}_{p_{2}}[T]-\operatorname{{\mathbb{E}}}_{p_{1}}[T]),$ and $\lambda^{*}$ is the unique scalar in $(0,1)$ such that the distribution $p^{*}\in\mathcal{P}$ of mean $\operatorname{{\mathbb{E}}}_{p^{*}}[T]=\tfrac{\lambda^{*}\operatorname{{\mathbb{E}}}_{p_{1}}[T]-\operatorname{{\mathbb{E}}}_{p_{2}}[T]}{\lambda^{*}-1}$ satisfies $\mathrm{KL}(p_{1}\|p^{*})=\alpha$.
  2. (ii)
    (Gaussian case) If $\mathcal{P}=\{\mathcal{N}(\mu,\Sigma):\mu\in\mathbb{R}^{d}\}$ with fixed $\Sigma\succ 0$, then for $\alpha\geq\mathrm{KL}(p_{1}\|p_{2})$:
    $v(\alpha)=\bigl{(}\sqrt{\alpha}-\sqrt{\mathrm{KL}(p_{1}\|p_{2})}\bigr{)}^{2}.$
Proof.
In this proof, for a fixed $\alpha\geq 0$, we study the optimal value
$v(\alpha)=\inf_{\theta\in\Theta}\bigl{\{}\mathrm{KL}(p_{2}\|p_{\theta})\;|\;\mathrm{KL}(p_{1}\|p_{\theta})\geq\alpha\bigr{\}}.$
This enables characterzing the Pareto frontier of feasible $(\alpha,\varepsilon)$ trade-offs. Below, we focus on the non-trivial case $\alpha>\mathrm{KL}(p_{1}\|p_{2})$. Indeed, in the case $\alpha\leq\mathrm{KL}(p_{1}\|p_{2})$, the problem above’s unconstrained minimizer $p_{2}$ is a feasible solution, so that $v(\alpha)=0$ for all $\alpha\leq\mathrm{KL}(p_{1}\|p_{2})$.
Reparametrization and KKT. First, we recall the expression of the KL divergence in exponential families as a Bregman divergence. For any $\theta^{\prime},\theta\in\Theta$ one has
$\mathrm{KL}(p_{\theta^{\prime}}\|p_{\theta})=A(\theta)-A(\theta^{\prime})-(\theta-\theta^{\prime})^{T}\nabla A(\theta^{\prime}).$
We let
$f(\theta)=\mathrm{KL}(p_{2}\|p_{\theta}),\quad g(\theta)=\mathrm{KL}(p_{1}\|p_{\theta}).$
Both are continuously differentiable on the open set $\Theta$. The feasible set $\{g(\theta)\geq\alpha\}$ is nonconvex, so we check linear independence constraint qualification (LICQ,  [7]) at any minimizer $\theta^{*}$. To do so, we recall (see [29]) that for exponential families $\nabla A(\theta)=\operatorname{{\mathbb{E}}}_{p_{\theta}}[T]$ for any $\theta\in\Theta$, so that
$\nabla g(\theta)=-\nabla A(\theta_{1})+\nabla A(\theta)=-\operatorname{{\mathbb{E}}}_{p_{1}}[T]+\operatorname{{\mathbb{E}}}_{p_{\theta}}[T].$
Minimality of the exponential family ensures $\operatorname{{\mathbb{E}}}_{p_{\theta}}[T]$ is injective [29]. This together with the fact that $p_{1}\neq p_{\theta}$ for any feasible $\theta$, since $\alpha>0$, implies that $\nabla g(\theta^{*})\neq 0$ and LICQ holds.
Next, we form the Lagrangian
$\mathcal{L}(\theta,\lambda)=f(\theta)-\lambda\bigl{(}g(\theta)-\alpha\bigr{)},\quad\lambda\geq 0.$
Since LICQ holds, KKT conditions are necessary [7] for any minimizer $\theta^{*}$. Hence, stationarity ($\nabla_{\theta}\mathcal{L}=0$) gives
$\nabla f(\theta^{*})=\lambda\,\nabla g(\theta^{*}).$
Given that we had derived $\nabla_{\theta}\mathrm{KL}(p_{i}\|p_{\theta})=-\operatorname{{\mathbb{E}}}_{p_{i}}[T]+\operatorname{{\mathbb{E}}}_{p_{\theta}}[T],$ we obtain
$-\operatorname{{\mathbb{E}}}_{p_{2}}[T]+\operatorname{{\mathbb{E}}}_{p^{*}}[T]=\lambda\,(-\operatorname{{\mathbb{E}}}_{p_{1}}[T]+\operatorname{{\mathbb{E}}}_{p^{*}}[T]),$
hence $(1-\lambda)\operatorname{{\mathbb{E}}}_{p^{*}}[T]=\operatorname{{\mathbb{E}}}_{p_{2}}[T]-\lambda\operatorname{{\mathbb{E}}}_{p_{1}}[T]$. Observe that we must have $\lambda\neq 1$, as otherwise $\operatorname{{\mathbb{E}}}_{p_{2}}[T]=\operatorname{{\mathbb{E}}}_{p_{1}}[T]$, but this contradicts the fact that $p_{1}\neq p_{1}$ given that $\operatorname{{\mathbb{E}}}_{p_{\theta}}[T]$ is injective for minimal exponential families [29]. Therefore, we have the following:
$\operatorname{{\mathbb{E}}}_{p^{*}}[T]=\frac{\lambda\operatorname{{\mathbb{E}}}_{p_{1}}[T]-\operatorname{{\mathbb{E}}}_{p_{2}}[T]}{\lambda-1}.$
Now, we observe that the inequality constraint must be active. Otherwise, the minimizer lies in the interior of the feasible set and is thus a local minimum of the unconstrained problem. The latter admits $p_{2}$ as a unique minimizer, so the minimizer at hand must be $p_{2}$ but this is not feasible since we assume $\mathrm{KL}(p_{1}\|p_{2})<\alpha$. Therefore, the minimizer must lie at the boundary of the feasible set, and the inequality constraint is active. That is, we have $g(\theta^{*})=\alpha$.
Optimal value. We are now ready to derive the expression of the optimal value:
$v(\alpha)=f(\theta^{\star})=\mathrm{KL}(p_{2}\|p^{*}).$
We use the following classical Pythagorean-type identity for Bregman divergences (see, e.g., Banerjee et al. [4]):
$\mathrm{KL}(p_{2}\|p^{*})=\mathrm{KL}(p_{2}\|p_{1})+\mathrm{KL}(p_{1}\|p^{*})+(\theta_{2}-\theta_{1})^{\top}(\nabla A(\theta_{1})-\nabla A(\theta^{*})).$
Now, consider the aforementioned KKT multiplier $\lambda>0,\lambda\neq 1$ such that $\nabla A(\theta^{*})=\operatorname{{\mathbb{E}}}_{p^{*}}[T]=\tfrac{\lambda\operatorname{{\mathbb{E}}}_{p_{1}}[T]-\operatorname{{\mathbb{E}}}_{p_{2}}[T]}{\lambda-1}=\tfrac{\lambda\nabla A(\theta_{1})-\nabla A(\theta_{2})}{\lambda-1}$ as well as $\mathrm{KL}(p_{1}\|p^{*})=g(\theta^{*})=\alpha$. We thus get
$v(\alpha)$ $=\mathrm{KL}(p_{2}\|p^{*})=\mathrm{KL}(p_{2}\|p_{1})+\mathrm{KL}(p_{1}\|p^{*})+(\theta_{2}-\theta_{1})^{\top}(\nabla A(\theta_{1})-\nabla A(\theta^{*}))$
$=\mathrm{KL}(p_{2}\|p_{1})+\mathrm{KL}(p_{1}\|p^{*})+\frac{1}{\lambda-1}(\theta_{2}-\theta_{1})^{\top}(\nabla A(\theta_{2})-\nabla A(\theta_{1}))$
$=\mathrm{KL}(p_{2}\|p_{1})+\alpha+\frac{1}{\lambda-1}(\theta_{2}-\theta_{1})^{\top}(\nabla A(\theta_{2})-\nabla A(\theta_{1})).$
We now again use that $\nabla A(\theta)=\operatorname{{\mathbb{E}}}_{p_{\theta}}[T]$ for all $\theta\in\Theta$. We then obtain:
$v(\alpha)$ $=\mathrm{KL}(p_{2}\|p_{1})+\alpha+\frac{1}{\lambda-1}(\theta_{2}-\theta_{1})^{\top}(\operatorname{{\mathbb{E}}}_{p_{2}}[T]-\operatorname{{\mathbb{E}}}_{p_{1}}[T]).$ (3)
Uniqueness of $\lambda$. We now show that there is a unique KKT multiplier $\lambda$ for the optimal solution. We recall that $\lambda>0$ is such that $\lambda\neq 1$ and:
$\operatorname{{\mathbb{E}}}_{p^{*}}[T]=\frac{\lambda\operatorname{{\mathbb{E}}}_{p_{1}}[T]-\operatorname{{\mathbb{E}}}_{p_{2}}[T]}{\lambda-1}$ $g(\theta^{*})=\mathrm{KL}(p_{1}\|p^{*})=\alpha.$
Above, the first equation uniquely defines the distribution $p^{*}$, by minimality of the family, and we now show that there is only a unique $\lambda$ of interest such that the second equations above holds.
Let $\theta^{*}(\lambda)$ be the unique (by minimality) parameter such that $\operatorname{{\mathbb{E}}}_{p^{*}}[T]=\tfrac{\lambda\operatorname{{\mathbb{E}}}_{p_{1}}[T]-\operatorname{{\mathbb{E}}}_{p_{2}}[T]}{\lambda-1}$. We define
$H(\lambda)\coloneqq\mathrm{KL}(p_{1}\|p^{*})=A(\theta^{*})-A(\theta_{1})-(\theta^{*}-\theta_{1})^{\top}\nabla A(\theta_{1})$
Taking the derivative above, we get
$\frac{dH}{d\lambda}(\lambda)=(\nabla A(\theta^{*})-\nabla A(\theta_{1}))^{\top}\frac{d\theta^{*}}{d\lambda}(\lambda).$
We again use that
$\nabla A(\theta^{*})=\operatorname{{\mathbb{E}}}_{p^{*}}[T]=\tfrac{\lambda\operatorname{{\mathbb{E}}}_{p_{1}}[T]-\operatorname{{\mathbb{E}}}_{p_{2}}[T]}{\lambda-1}=\frac{\lambda\nabla A(\theta_{1})-\nabla A(\theta_{2})}{\lambda-1}.$
First, replacing in the previous derivative equation, we obtain:
$\frac{dH}{d\lambda}(\lambda)=(\nabla A(\theta^{*})-\nabla A(\theta_{1}))^{\top}\frac{d\theta^{*}}{d\lambda}(\lambda)=\frac{1}{\lambda-1}(\nabla A(\theta_{1})-\nabla A(\theta_{2}))^{\top}\frac{d\theta^{*}}{d\lambda}(\lambda).$
Second, taking the derivative with respect to $\lambda$ in the expression of $\nabla A(\theta^{*})$ yields:
$\nabla^{2}A(\theta^{*})\cdot\frac{d\theta^{*}}{d\lambda}(\lambda)=\frac{1}{(\lambda-1)^{2}}(\nabla A(\theta_{1})-\nabla A(\theta_{2})).$
We observe that the Fisher information matrix $\nabla^{2}A(\theta^{*})$ is positive definite since the family is regular. Multiplying by the inverse of the latter then yields:
$\frac{d\theta^{*}}{d\lambda}(\lambda)=\frac{1}{(\lambda-1)^{2}}\nabla^{2}A(\theta^{*})^{-1}(\nabla A(\theta_{1})-\nabla A(\theta_{2})).$
Plugging the above in the latest expression of the derivative of $H$ yields:
$\frac{dH}{d\lambda}(\lambda)$ $=\frac{1}{\lambda-1}(\nabla A(\theta_{1})-\nabla A(\theta_{2}))^{\top}\frac{d\theta^{*}}{d\lambda}(\lambda)$
$=\frac{1}{(\lambda-1)^{3}}(\nabla A(\theta_{1})-\nabla A(\theta_{2}))^{\top}\nabla^{2}A(\theta^{*})^{-1}(\nabla A(\theta_{1})-\nabla A(\theta_{2})).$
Since the matrix $\nabla^{2}A(\theta^{*})$ is positive definite by regularity of the family, the corresponding quadratic form above is positive (recall $\theta_{1}\neq\theta_{2}$), and the sign of the derivative is that of $\lambda-1$. Therefore, $H$ is decreasing on $(0,1)$ and increasing on $(1,+\infty)$. It is straighforward to check that $H(0^{+})=\mathrm{KL}(p_{1}\|p_{2})$, $H(1^{-})=H(1^{+})=+\infty$, and $H(+\infty)=0$. Since $\mathrm{KL}(p_{1}\|p_{2})<\alpha$ by assumption, there exists a unique $\lambda^{*}\in(0,1)$ such that $H(\lambda^{*})=\alpha$ and a unique $\lambda^{*}_{1}>1$ such that $H(\lambda^{*}_{1})=\alpha$.
We now discard $\lambda^{*}_{1}$ thanks to the expression of the optimal value expression (3). Indeed, the second term in (3) is positive for $\lambda^{*}_{1}$ since $\lambda^{*}_{1}>1$ and $(\theta_{2}-\theta_{1})^{\top}(\operatorname{{\mathbb{E}}}_{p_{2}}[T]-\operatorname{{\mathbb{E}}}_{p_{1}}[T])=(\theta_{2}-\theta_{1})^{\top}(\nabla A(\theta_{2})-\nabla A(\theta_{1}))>0$ by strict convexity of $A$ and the fact that $p_{1}\neq p_{2}$. On the other hand, this same second term is negative for $\lambda_{*}$ since $\lambda_{*}<1$. Therefore, the optimal value is smaller for the choice of $\lambda^{*}$, so that:
$v(\alpha)$ $=\mathrm{KL}(p_{2}\|p_{1})+\alpha+\frac{1}{\lambda^{*}-1}(\theta_{2}-\theta_{1})^{\top}(\operatorname{{\mathbb{E}}}_{p_{2}}[T]-\operatorname{{\mathbb{E}}}_{p_{1}}[T]),$ (4)
where $\lambda^{*}$ is the unique scalar in $(0,1)$ such that:
$\operatorname{{\mathbb{E}}}_{p^{*}}[T]=\frac{\lambda^{*}\operatorname{{\mathbb{E}}}_{p_{1}}[T]-\operatorname{{\mathbb{E}}}_{p_{2}}[T]}{\lambda^{*}-1}$ $g(\theta^{*})=\mathrm{KL}(p_{1}\|p^{*})=\alpha.$
Finally, we note that $v(\alpha)$ is non-decreasing by definition; increasing $\alpha$ shrinks the feasible set. Also, we recall that $v(\alpha)=0$ for all $\alpha<\mathrm{KL}(p_{1}\|p_{2})$, i.e., all trade-offs $(\alpha,\varepsilon)$ with $\alpha<\mathrm{KL}(p_{1}\|p_{2}),\varepsilon>0$ are dominated by $(\mathrm{KL}(p_{1}\|p_{2}),0)$. Therefore, we conclude that the pareto frontier is given by:
$\mathrm{PF}(p_{1},p_{2};\mathcal{P})=\Bigl{\{}\bigl{(}\alpha,\,v(\alpha)\bigr{)}:\alpha\geq\mathrm{KL}(p_{1}\|p_{2})\Bigr{\}}.$
Gaussian case. For $p_{\mu}=\mathcal{N}(\mu,\Sigma)$ one has $T(x)=x$, $E_{p_{\mu}}[T]=\mu$, and
$\mathrm{KL}(p_{i}\|p_{\mu})=\tfrac{1}{2}(\mu_{i}-\mu)^{\top}\Sigma^{-1}(\mu_{i}-\mu).$
By the conditions on $\lambda^{*}$ we have
$\mu^{*}=\frac{\lambda^{*}\mu_{1}-\mu_{2}}{\lambda^{*}-1},\quad\mathrm{KL}(p_{1}\|p_{\mu^{*}})=\alpha.$
Using the KL divergence expression for Gaussians $\mathcal{N}(\mu,\Sigma)$ (recall $\mathrm{KL}(\mathcal{N}(\mu_{1},\Sigma),\mathcal{N}(\mu_{2},\Sigma))=\tfrac{1}{2}(\mu_{1}-\mu_{2})^{\top}\Sigma^{-1}(\mu_{1}-\mu_{2})$), we get
$\mu_{1}-\mu^{*}=\frac{\mu_{2}-\mu_{1}}{\lambda^{*}-1},\quad\alpha=\frac{\mathrm{KL}(p_{1}\|p_{2})}{(\lambda^{*}-1)^{2}}.$
Solving for $\lambda^{*}\in(0,1)$ yields $\lambda^{*}=1-\sqrt{\mathrm{KL}(p_{1}\|p_{2})/\alpha}$ and
$\mu^{*}=\mu_{1}+\sqrt{\frac{\alpha}{\mathrm{KL}(p_{1}\|p_{2})}}(\mu_{2}-\mu_{1}).$
Thus, direct computations yield
$v(\alpha)=\tfrac{1}{2}\Bigl{(}\|\mu_{2}-\mu_{1}\|_{\Sigma^{-1}}-\sqrt{2\alpha}\Bigr{)}^{2}=\bigl{(}\sqrt{\mathrm{KL}(p_{1}\|p_{2})}-\sqrt{\alpha}\bigr{)}^{2},$
with $v(\alpha)=0$ if $\alpha\leq\mathrm{KL}(p_{1}\|p_{2})$. This concludes the proof. ∎
Discussion.
In Proposition 3 we show that, in any regular exponential family, the trade-off between removal ($\alpha$) and preservation ($\varepsilon$) can be quantified. This yields a removal-preservation trade-off curve that faithfully reproduces the shared-covariance Gaussian Pareto frontier—namely the familiar $(\sqrt{\alpha}-\sqrt{D})^{2}$ parabola—while in other families it gives an explicit but generally non-algebraic trade-off curve.
A.2 Predictive Performance
See 2
Proof.
Let $h(y\mid x):=h(x)(y)$ denote the conditional distribution defined by the hypothesis $h$, and suppose $h$ minimizes the expected log-loss under $p$. Since $\ell(y,q)=-\log q(y)$ is a strictly proper scoring rule, the unique minimizer of $\mathcal{L}(h;p)$ is the true conditional distribution $h(x)=p(\cdot\mid x)$, where $p(x,y)=p^{X}(x)p(y\mid x)$.
We begin by analyzing the expected log-loss of this hypothesis under an arbitrary distribution $q$ over $\mathcal{X}\times\mathcal{Y}$:
$\mathcal{L}(h;q)$ $=\mathbb{E}_{(x,y)\sim q}[-\log h(y\mid x)]=\mathbb{E}_{x\sim q^{X}}\mathbb{E}_{y\sim q(\cdot\mid x)}[-\log p(y\mid x)],$
where $q^{X}$ denotes the marginal distribution of $x$ under $q$, and $q(\cdot\mid x)$ the corresponding conditional.
Now recall the standard identity for any two conditional distributions $q(\cdot\mid x)$ and $p(\cdot\mid x)$:
$\mathbb{E}_{y\sim q(\cdot\mid x)}[-\log p(y\mid x)]=\mathrm{KL}(q(y\mid x)\parallel p(y\mid x))+H(q(y\mid x)),$
where $H(q(y\mid x))=\mathbb{E}_{y\sim q(\cdot\mid x)}[-\log q(y\mid x)]$ is the Shannon entropy of the label distribution under $q$ for fixed $x$.
Taking the expectation over $x\sim q^{X}$, we get:
$\mathcal{L}(h;q)$ $=\mathbb{E}_{x\sim q^{X}}\left[\mathrm{KL}(q(y\mid x)\parallel p(y\mid x))+H(q(y\mid x))\right]$
$=\mathbb{E}_{x\sim q^{X}}\mathrm{KL}(q(y\mid x)\parallel p(y\mid x))+\mathbb{E}_{x\sim q^{X}}H(q(y\mid x)).$
The second term is the expected entropy, which corresponds to the Bayes-optimal risk under $q$:
$\mathcal{L}(h_{q}^{*};q):=\inf_{h^{\prime}}\mathcal{L}(h^{\prime};q)=\mathbb{E}_{x\sim q^{X}}H(q(y\mid x)).$ (5)
Next, we relate the expected conditional KL term to the total KL divergence between the joint distributions. Using the chain rule for KL divergence, we have:
$\mathrm{KL}(q\parallel p)=\mathrm{KL}(q^{X}\parallel p^{X})+\mathbb{E}_{x\sim q^{X}}\mathrm{KL}(q(y\mid x)\parallel p(y\mid x)).$ (6)
This decomposition holds generally for joint distributions with conditional factorizations. Solving for the conditional KL term, we obtain:
$\mathbb{E}_{x\sim q^{X}}\mathrm{KL}(q(y\mid x)\parallel p(y\mid x))=\mathrm{KL}(q\parallel p)-\mathrm{KL}(q^{X}\parallel p^{X}).$ (7)
Substituting the above into the expression for $\mathcal{L}(h;q)$ and using (5), we get:
$\mathcal{L}(h;q)=\mathrm{KL}(q\parallel p)-\mathrm{KL}(q^{X}\parallel p^{X})+\mathcal{L}(h_{q}^{*};q).$ (8)
We now apply this to $q=p_{1}$ and $q=p_{2}$, noting that $p$ satisfies $(\alpha,\varepsilon)$-distributional unlearning, i.e., $\mathrm{KL}(p_{1}\parallel p)\geq\alpha$ and $\mathrm{KL}(p_{2}\parallel p)\leq\varepsilon$.
For $p_{1}$, we define $\delta_{1}:=\mathrm{KL}(p_{1}^{X}\parallel p^{X})$ and compute:
$\mathcal{L}(h;p_{1})-\mathcal{L}(h_{1};p_{1})$ $=\mathrm{KL}(p_{1}\parallel p)-\mathrm{KL}(p_{1}^{X}\parallel p^{X})=\mathrm{KL}(p_{1}\parallel p)-\delta_{1}\geq\alpha-\delta_{1}.$
For $p_{2}$, define $\delta_{2}:=\mathrm{KL}(p_{2}^{X}\parallel p^{X})$ and similarly compute:
$\mathcal{L}(h;p_{2})-\mathcal{L}(h_{2};p_{2})$ $=\mathrm{KL}(p_{2}\parallel p)-\mathrm{KL}(p_{2}^{X}\parallel p^{X})=\mathrm{KL}(p_{2}\parallel p)-\delta_{2}\leq\varepsilon-\delta_{2}.$
This completes the proof. ∎
A.3 Random Removal
Lemma 3 (Finite-sample concentration).
Let $\hat{\mu}$ be the empirical mean of $n$ samples drawn from $\mathcal{N}(\mu,\sigma^{2})$. For any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
$|\hat{\mu}-\mu|\leq\sigma\sqrt{\frac{2\ln(2/\delta)}{n}}.$
Proof.
This follows directly from Hoeffding’s inequality for sub-Gaussian variables. ∎
See 1
Proof.
We recall that $p_{1}=\mathcal{N}(\mu_{1},\sigma^{2}),p_{2}=\mathcal{N}(\mu_{2},\sigma^{2})\in\mathcal{P}$ and $p\coloneqq\mathcal{N}(\mu,\sigma^{2})\in\mathcal{P}$ are univariate Gaussian distributions. We are given $n_{1}$ i.i.d. samples $x_{1}^{(1)},\ldots,x_{1}^{(n_{1})}$ from $p_{1}$ and $n_{2}$ i.i.d. samples $x_{2}^{(1)},\ldots,x_{2}^{(n_{2})}$ from $p_{2}$.
Upon removing $f\leq n_{1}$ randomly chosen samples $x_{1}^{(1)},\ldots,x_{1}^{(n_{1}-f)}$ from the target distribution $p_{1}$, we set the center $\mu$ of the unlearned distribution $p$ to be:
$\mu=\frac{(n_{1}-f)\hat{\mu}_{1}+n_{2}\hat{\mu}_{2}}{n_{1}-f+n_{2}},$ (9)
where $\hat{\mu}_{1}\coloneqq\tfrac{1}{n_{1}-f}\sum_{i=1}^{n_{1}-f}x_{1}^{(i)}$ and $\hat{\mu}_{2}\coloneqq\tfrac{1}{n_{2}}\sum_{i=1}^{n_{2}}x_{2}^{(i)}$. We also observe that a standard Hoeffding bound (Lemma 3) yields that:
$\left|\hat{\mu}_{1}-\mu_{1}\right|\leq\sigma\sqrt{\frac{2\ln(4/\delta)}{f}},\quad\left|\hat{\mu}_{2}-\mu_{2}\right|\leq\sigma\sqrt{\frac{2\ln(4/\delta)}{n_{2}}},$ (10)
each with probability $1-\tfrac{\delta}{2}$, so that both hold with probability $1-\delta$ thanks to a union bound. We also recall that
$\mathrm{KL}(p_{1}\parallel p)=\frac{(\mu_{1}-\mu)^{2}}{2\sigma^{2}},\quad\mathrm{KL}(p_{2}\parallel p)=\frac{(\mu_{2}-\mu)^{2}}{2\sigma^{2}}.$ (11)
Preservation bound.
First, we upper bound the KL divergence of $p_{2}$ from $p$. To do so, we first use the triangle inequality to get
$|\mu-\mu_{2}|$ $=\left|\frac{(n_{1}-f)\hat{\mu}_{1}+n_{2}\hat{\mu}_{2}}{n_{1}-f+n_{2}}-\mu_{2}\right|=\left|\frac{n_{1}-f}{n_{1}-f+n_{2}}(\hat{\mu}_{1}-\mu_{2})+\frac{n_{2}}{n_{1}-f+n_{2}}(\hat{\mu}_{2}-\mu_{2})\right|$
$=\left|\frac{n_{1}-f}{n_{1}-f+n_{2}}(\mu_{1}-\mu_{2})+\frac{n_{1}-f}{n_{1}-f+n_{2}}(\hat{\mu}_{1}-\mu_{1})+\frac{n_{2}}{n_{1}-f+n_{2}}(\hat{\mu}_{2}-\mu_{2})\right|$
$\leq\frac{n_{1}-f}{n_{1}-f+n_{2}}\left|\mu_{1}-\mu_{2}\right|+\frac{n_{1}-f}{n_{1}-f+n_{2}}\left|\hat{\mu}_{1}-\mu_{1}\right|+\frac{n_{2}}{n_{1}-f+n_{2}}\left|\hat{\mu}_{2}-\mu_{2}\right|.$
Therefore, using (10) we have with probability $1-\delta$:
$|\mu-\mu_{2}|$ $\leq\frac{n_{1}-f}{n_{1}-f+n_{2}}\left|\mu_{1}-\mu_{2}\right|+\frac{n_{1}-f}{n_{1}-f+n_{2}}\sigma\sqrt{\frac{2\ln(4/\delta)}{f}}+\frac{n_{2}}{n_{1}-f+n_{2}}\sigma\sqrt{\frac{2\ln(4/\delta)}{n_{2}}}.$
Taking squares, using Jensen’s inequality, and simplifying further since $f\geq 0$, yields:
$|\mu-\mu_{2}|^{2}$ $\leq 3\left(\frac{n_{1}-f}{n_{2}}\right)^{2}\left|\mu_{1}-\mu_{2}\right|^{2}+3\left(\frac{n_{1}-f}{n_{2}}\right)^{2}\sigma^{2}\frac{2\ln(4/\delta)}{n_{1}-f}+\sigma^{2}\frac{6\ln(4/\delta)}{n_{2}}.$ (12)
Dividing both sides by $2\sigma^{2}$ and then using (28) yields with probability $1-\delta$:
$\mathrm{KL}(p_{2}\parallel p)$ $\leq 3\left(\frac{n_{1}-f}{n_{2}}\right)^{2}\mathrm{KL}(p_{1}\parallel p_{2})+\frac{3\ln(4/\delta)}{n_{2}}\left(1+\frac{n_{1}-f}{n_{2}}\right).$ (13)
Removal bound.
Second, we lower bound the KL divergence of $p_{1}$ from $p$. To do so, we use Jensen’s inequality and (12) to obtain that, with probability $1-\delta$, we have
$|\mu_{1}-\mu_{2}|^{2}$ $=|\mu_{1}-\mu+\mu-\mu_{2}|^{2}\leq 2|\mu_{1}-\mu|^{2}+2|\mu-\mu_{2}|^{2}$
$\leq 2|\mu_{1}-\mu|^{2}+6\left(\frac{n_{1}-f}{n_{2}}\right)^{2}\left|\mu_{1}-\mu_{2}\right|^{2}+\frac{6\sigma^{2}\ln(4/\delta)}{n_{2}}\left(1+\frac{n_{1}-f}{n_{2}}\right)$
Rearranging terms and dividing by $4\sigma^{2}$ along with (28) yields that with probability $1-\delta$ we have
$\mathrm{KL}(p_{1}\parallel p)$ $=\frac{|\mu_{1}-\mu|^{2}}{2\sigma^{2}}\geq\frac{|\mu_{1}-\mu_{2}|^{2}}{4\sigma^{2}}-3\left(\frac{n_{1}-f}{n_{2}}\right)^{2}\frac{\left|\mu_{1}-\mu_{2}\right|^{2}}{2\sigma^{2}}-\frac{3\ln(4/\delta)}{2n_{2}}\left(1+\frac{n_{1}-f}{n_{2}}\right)$
$=\left(\frac{1}{2}-3\left(\frac{n_{1}-f}{n_{2}}\right)^{2}\right)\mathrm{KL}(p_{1}\parallel p_{2})-\frac{3\ln(4/\delta)}{2n_{2}}\left(1+\frac{n_{1}-f}{n_{2}}\right).$
Conclusion.
With probability $1-\delta$, we have $(\alpha,\varepsilon)$-distributional unlearning with
$\alpha$ $\geq\left(\frac{1}{2}-3\left(\frac{n_{1}-f}{n_{2}}\right)^{2}\right)\mathrm{KL}(p_{1}\parallel p_{2})-\frac{3\ln(4/\delta)}{2n_{2}}\left(1+\frac{n_{1}-f}{n_{2}}\right),$
$\varepsilon$ $\leq 3\left(\frac{n_{1}-f}{n_{2}}\right)^{2}\mathrm{KL}(p_{1}\parallel p_{2})+\frac{3\ln(4/\delta)}{n_{2}}\left(1+\frac{n_{1}-f}{n_{2}}\right).$
A.4 Selective Removal
Lemma 4 (Dvoretzky–Kiefer–Wolfowitz Inequality).
Let $x_{1},x_{2},\dots,x_{n}$ be independent and identically distributed random variables with cumulative distribution function $F$. Define the empirical distribution function by
$\widehat{F}(t)=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{x_{i}\leq t\}.$
Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$ we have
$\sup_{t\in\mathbb{R}}\left|\widehat{F}(t)-F(t)\right|\leq\sqrt{\frac{\ln(2/\delta)}{2n}}.$
Lemma 5.
Let $\mu_{1},\mu_{2}\in\mathbb{R}$, $\sigma>0$. Consider $n_{1}$ i.i.d. samples $x_{1}^{(1)},\ldots,x_{1}^{(n_{1})}$ from $\mathcal{N}(\mu_{1},\sigma^{2})$ and $n_{2}$ i.i.d. samples $x_{2}^{(1)},\ldots,x_{2}^{(n_{2})}$ from $\mathcal{N}(\mu_{2},\sigma^{2})$. We define $\hat{\mu}_{2}$ the average of the samples from $\mathcal{N}(\mu_{2},\sigma^{2})$, $\hat{\mu}_{1}$ the average of the $n_{1}-f\leq n_{1}$ closest samples from $x_{1}^{(1)},\ldots,x_{1}^{(n_{1})}$ to $\hat{\mu}_{2}$. We define $F:t\in\mathbb{R}\mapsto\Phi(\frac{t-|\mu_{1}-\mu_{2}|}{\sigma})-\Phi(\frac{-t-|\mu_{1}-\mu_{2}|}{\sigma})$, where $\Phi$ is the standard normal CDF.
For any $\delta\in(0,1)$, we have with probability $1-\delta$,
$\left|\hat{\mu}_{1}-\mu_{2}\right|\leq F^{-1}\left(1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(2/\delta)}{2n_{1}}}\right).$ (14)
Proof.
Recall from Equation (25) that $\hat{\mu}_{1}$ is the average of the $n_{1}-f$ samples, out of $n_{1}$ i.i.d. from $p_{1}$, with the closest distance to $\hat{\mu}_{2}$, the empirical mean of $n_{2}$ samples from $p_{2}$.
Denote by $\hat{\tau}_{f}\coloneqq|x_{1}^{(n_{1}-f:n_{1})}-\hat{\mu}_{2}|$ the $(n_{1}-f)$-th largest distance of $\hat{\mu}_{2}$ to $p_{1}$ samples. It is then immediate from the triangle inequality that
$|\hat{\mu}_{1}-\mu_{2}|=|\frac{1}{n_{1}-f}\sum_{i=1}^{n_{1}-f}x_{1}^{(i:n_{1})}-\mu_{2}|\leq\frac{1}{n_{1}-f}\sum_{i=1}^{n_{1}-f}|x_{1}^{(i:n_{1})}-\mu_{2}|\leq\hat{\tau}_{f}.$ (15)
Besides, denoting by $\widehat{F}_{1}$ the empirical CDF of the empirical distribution over $\left\{|x_{1}^{(i)}-\mu_{2}|~{}\colon~{}i\in[n_{1}]\right\}$, we have for all $t\in\mathbb{R}$:
$\widehat{F}_{1}(t)=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\mathbbm{1}_{\left\{|x_{1}^{(i)}-\mu_{2}|\leq t\right\}}.$ (16)
Yet, we recall that with probability $1-\tfrac{\delta}{2}$, we have
$\left|\hat{\mu}_{2}-\mu_{2}\right|\leq\sigma\sqrt{\frac{2\ln(4/\delta)}{n_{2}}}.$ (17)
Therefore, the triangle inequality gives for $i\in[n_{1}]$, $|x_{1}^{(i)}-\mu_{2}|\leq|x_{1}^{(i)}-\hat{\mu}_{2}|+|\hat{\mu}_{2}-\mu|\leq|x_{1}^{(i)}-\hat{\mu}_{2}|+\sigma\sqrt{\tfrac{2\ln(4/\delta)}{n_{2}}}$, and we deduce
$\widehat{F}_{1}(t)=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\mathbbm{1}_{\left\{|x_{1}^{(i)}-\mu_{2}|\leq t\right\}}\leq\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\mathbbm{1}_{\left\{|x_{1}^{(i)}-\hat{\mu}_{2}|\leq t-\sigma\sqrt{\tfrac{2\ln(4/\delta)}{n_{2}}}\right\}}.$ (18)
In particular, by definition of $\hat{\tau}_{f}$, we have
$\widehat{F}_{1}(\hat{\tau}_{f}+\sigma\sqrt{\frac{2\ln(4/\delta)}{n_{2}}})\leq\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\mathbbm{1}_{\left\{|x_{1}^{(i)}-\hat{\mu}_{2}|\leq\hat{\tau}_{f}\right\}}=\frac{n_{1}-f}{n_{1}}=1-\frac{f}{n_{1}}.$ (19)
Now, observe that $|x_{1}^{(i)}-\mu_{2}|$ follows a folded normal distribution of location $\mu_{1}-\mu_{2}$ and scale $\sigma^{2}$, since $x_{1}^{(i)}$ follows $p_{1}=\mathcal{N}(\mu_{1},\sigma^{2})$. Denote by $F_{1}$ its CDF. Thanks to the Dvoretzky–Kiefer–Wolfowitz inequality (Lemma 4), we have with probability $1-\tfrac{\delta}{2}$ that for all $t\in\mathbb{R}$,
$|\widehat{F}_{1}(t)-F_{1}(t)|\leq\sqrt{\frac{\ln(4/\delta)}{2n_{1}}}.$ (20)
Plugging the above in the previous inequality, and using a union bound, we get with probability $1-\delta$,
$F_{1}(\hat{\tau}_{f}+\sigma\sqrt{\frac{2\ln(4/\delta)}{n_{2}}})\leq\widehat{F}_{1}(\hat{\tau}_{f}+\sigma\sqrt{\frac{2\ln(4/\delta)}{n_{2}}})+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}}\leq 1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}}.$ (21)
By taking the inverse $F^{-1}_{1}$ of the CDF $F_{1}$ and rearranging terms, we obtain with probability $1-\delta$ that
$\hat{\tau}_{f}\leq F_{1}^{-1}{\left(1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}}\right)}-\sigma\sqrt{\frac{2\ln(4/\delta)}{n_{2}}}.$ (22)
Finally, going back to (15), we obtain with probability $1-\delta$ that
$|\hat{\mu}_{1}-\mu_{2}|\leq\hat{\tau}_{f}\leq F_{1}^{-1}{\left(1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}}\right)}-\sigma\sqrt{\frac{2\ln(4/\delta)}{n_{2}}}.$ (23)
Lemma 6.
Let $\mu_{1},\mu_{2}\in\mathbb{R}$ and suppose that $x\sim\mathcal{N}(\mu_{1},\sigma^{2})$. Define the random variable $z\coloneqq|x-\mu_{2}|,$ with cumulative distribution function
$F_{\sigma}(t)\coloneqq\mathbb{P}[z\leq t]=\Phi\Bigl{(}\frac{t-|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}-\Phi\Bigl{(}\frac{-t-|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)},\quad t\geq 0,$
where $\Phi$ denotes the standard normal CDF. Then, for any $p\in(0,1)$ the inverse CDF satisfies
$F_{\sigma}^{-1}(p)=\sigma\,g^{-1}\Bigl{(}p;\frac{|\mu_{1}-\mu_{2}|^{2}}{2\sigma^{2}}\Bigr{)},$
where the function $g(u;\kappa)$ is defined by $g(u;\kappa)\coloneqq\Phi(u-\sqrt{2\kappa})+\Phi(u+\sqrt{2\kappa})-1,$ and $g^{-1}(p;\kappa)$ denotes the inverse function in $u$ satisfying $g(u;\kappa)=p$. In particular, when $\mu_{1}=\mu_{2}$ (so that $\kappa=0$) we have $g(u;0)=2\Phi(u)-1$ and thus $F_{0,1}^{-1}(p)=\Phi^{-1}\Bigl{(}\frac{p+1}{2}\Bigr{)}.$
Proof.
Since $x\sim\mathcal{N}(\mu_{1},\sigma^{2})$, we have that $z=|x-\mu_{2}|$ has CDF
$F_{\sigma}(t)=\Phi\Bigl{(}\frac{t-|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}-\Phi\Bigl{(}\frac{-t-|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)},\quad t\geq 0.$
Introduce the change of variable $u=\frac{t}{\sigma}$ so that $t=\sigma\,u$. Then,
$F_{\sigma}(\sigma\,u)=\Phi\Bigl{(}u-\frac{|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}-\Phi\Bigl{(}-u-\frac{|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}.$
Using the symmetry $\Phi(-x)=1-\Phi(x)$, this becomes
$F_{\sigma}(\sigma\,u)=\Phi\Bigl{(}u-\frac{|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}+\Phi\Bigl{(}u+\frac{|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}-1.$
Defining $\kappa=\frac{|\mu_{1}-\mu_{2}|^{2}}{2\sigma^{2}}$ and setting
$g(u;\kappa)\coloneqq\Phi(u-\sqrt{2\kappa})+\Phi(u+\sqrt{2\kappa})-1,$
we have $F_{\sigma}(\sigma\,u)=g(u;\kappa)$. Thus, if $u^{*}$ is the unique solution of $g(u^{*};\kappa)=p$, then
$F_{\sigma}(\sigma\,u^{*})=p,$
so that
$F_{\sigma}^{-1}(p)=\sigma\,u^{*}=\sigma\,g^{-1}(p;\kappa).$
In the special case $\mu_{1}=\mu_{2}$ (so that $\kappa=0$), we obtain $g(u;0)=2\Phi(u)-1$, whose inverse is given by $u=\Phi^{-1}((p+1)/2)$. Hence, $F_{1}^{-1}(p)=\Phi^{-1}((p+1)/2)$, as required. ∎
See 2
Proof.
We recall that $p_{1}=\mathcal{N}(\mu_{1},\sigma^{2}),p_{2}=\mathcal{N}(\mu_{2},\sigma^{2})\in\mathcal{P}$ and $p\coloneqq\mathcal{N}(\mu,\sigma^{2})\in\mathcal{P}$ are univariate Gaussian distributions. We are given $n_{1}$ i.i.d. samples $x_{1}^{(1)},\ldots,x_{1}^{(n_{1})}$ from $p_{1}$ and $n_{2}$ i.i.d. samples $x_{2}^{(1)},\ldots,x_{2}^{(n_{2})}$ from $p_{2}$.
The distance-based selection removes $f\leq n_{1}$ selected samples from the target distribution $p_{1}$ with the $f$ largest distances to $\hat{\mu}_{2}\coloneqq\tfrac{1}{n_{2}}\sum_{i=1}^{n_{2}}x_{2}^{(i)}$ the empirical estimator of the mean of $p_{2}$. That is, denoting by $x_{1}^{(1:n_{1})},\ldots,x_{1}^{(n_{1}:n_{1})}$ the original $n_{1}$ samples from $p_{1}$ reordered by increasing distance to $\hat{\mu}_{2}$:
$|x_{1}^{(1:n_{1})}-\hat{\mu}_{2}|\leq\ldots\leq|x_{1}^{(n_{1}:n_{1})}-\hat{\mu}_{2}|,$ (24)
with ties broken arbitrarily, then distance-based selection retains only $x_{1}^{(1:n_{1})},\ldots,x_{1}^{(n_{1}-f:n1)}$ to obtain
$\hat{\mu}_{1}\coloneqq\tfrac{1}{n_{1}-f}\sum_{i=1}^{n_{1}-f}x_{1}^{(i:n_{1})}.$ (25)
Subsequently, we set the center $\mu$ of the unlearned distribution $p$ to be:
$\mu=\frac{(n_{1}-f)\hat{\mu}_{1}+n_{2}\hat{\mu}_{2}}{n_{1}-f+n_{2}},$ (26)
where $\hat{\mu}_{1}=\tfrac{1}{n_{1}-f}\sum_{i=1}^{n_{1}-f}x_{1}^{(i:n_{1})}$ and $\hat{\mu}_{2}=\tfrac{1}{n_{2}}\sum_{i=1}^{n_{2}}x_{2}^{(i)}$. We also observe that a standard Hoeffding bound (Lemma 3) yields that:
$\left|\hat{\mu}_{2}-\mu_{2}\right|\leq\sigma\sqrt{\frac{2\ln(4/\delta)}{n_{2}}},$ (27)
with probability $1-\tfrac{\delta}{2}$. We also recall that
$\mathrm{KL}(p_{1}\parallel p)=\frac{(\mu_{1}-\mu)^{2}}{2\sigma^{2}},\quad\mathrm{KL}(p_{2}\parallel p)=\frac{(\mu_{2}-\mu)^{2}}{2\sigma^{2}}.$ (28)
Preservation bound.
First, we upper bound the KL divergence of $p_{2}$ from $p$. To do so, we first use the triangle inequality to get
$|\mu-\mu_{2}|$ $=\left|\frac{(n_{1}-f)\hat{\mu}_{1}+n_{2}\hat{\mu}_{2}}{n_{1}-f+n_{2}}-\mu_{2}\right|=\left|\frac{n_{1}-f}{n_{1}-f+n_{2}}(\hat{\mu}_{1}-\mu_{2})+\frac{n_{2}}{n_{1}-f+n_{2}}(\hat{\mu}_{2}-\mu_{2})\right|$
$\leq\frac{n_{1}-f}{n_{1}-f+n_{2}}\left|\hat{\mu}_{1}-\mu_{2}\right|+\frac{n_{2}}{n_{1}-f+n_{2}}\left|\hat{\mu}_{2}-\mu_{2}\right|.$
Therefore, using (27) we have with probability $1-\tfrac{\delta}{2}$:
$|\mu-\mu_{2}|$ $\leq\frac{n_{1}-f}{n_{1}-f+n_{2}}\left|\hat{\mu}_{1}-\mu_{2}\right|+\frac{n_{2}}{n_{1}-f+n_{2}}\sigma\sqrt{\frac{2\ln(4/\delta)}{n_{2}}}.$
Moreover, we know from Lemma 5 that with probabilty $1-\tfrac{\delta}{2}$
$\left|\hat{\mu}_{1}-\mu_{2}\right|\leq F^{-1}\left(1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}}\right).$
Using the above in the previous inequality with a union bound, yields that with probability $1-\delta$
$|\mu-\mu_{2}|$ $\leq\frac{n_{1}-f}{n_{1}-f+n_{2}}F^{-1}\left(1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}}\right)+\frac{n_{2}}{n_{1}-f+n_{2}}\sigma\sqrt{\frac{2\ln(4/\delta)}{n_{2}}}.$
We can further simplify the above using Lemma 6, which implies that for all $p>0$
$F^{-1}(p)=\sigma\,g^{-1}\Bigl{(}p;\frac{|\mu_{1}-\mu_{2}|^{2}}{2\sigma^{2}}\Bigr{)},$
where the function $g$ is defined by $g(u;\kappa)\coloneqq\Phi(u-\sqrt{2\kappa})+\Phi(u+\sqrt{2\kappa})-1$, for $u,\kappa>0$. Plugging this in the previous bound yields
$|\mu-\mu_{2}|$ $\leq\frac{n_{1}-f}{n_{1}-f+n_{2}}\sigma\,g^{-1}\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}};\frac{|\mu_{1}-\mu_{2}|^{2}}{2\sigma^{2}}\Bigr{)}+\frac{n_{2}}{n_{1}-f+n_{2}}\sigma\sqrt{\frac{2\ln(4/\delta)}{n_{2}}}.$ (29)
Dividing both sides by $\sigma\sqrt{2}$ and then using (28) yields with probability $1-\delta$:
$\sqrt{\mathrm{KL}(p_{2}\parallel p)}$ $\leq\frac{n_{1}-f}{(n_{1}-f+n_{2})\sqrt{2}}g^{-1}\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}};\mathrm{KL}(p_{1}\parallel p_{2})\Bigr{)}+\frac{\sqrt{n_{2}\log(4/\delta)}}{n_{1}-f+n_{2}}.$ (30)
The above directly implies, by taking squares and using Jensen’s inequality and that $f\geq 0$, that with probability $1-\delta$:
$\mathrm{KL}(p_{2}\parallel p)$ $\leq\left(\frac{n_{1}-f}{n_{2}}\right)^{2}g^{-1}\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}};\mathrm{KL}(p_{1}\parallel p_{2})\Bigr{)}^{2}+\frac{2\ln(4/\delta)}{n_{2}}.$ (31)
Removal bound.
Second, we lower bound the KL divergence of $p_{1}$ from $p$. To do so, we use Jensen’s inequality and (29) to obtain that, with probability $1-\delta$, we have
$|\mu_{1}-\mu_{2}|^{2}$ $=|\mu_{1}-\mu+\mu-\mu_{2}|^{2}\leq 2|\mu_{1}-\mu|^{2}+2|\mu-\mu_{2}|^{2}$
$\leq 2|\mu_{1}-\mu|^{2}+2\left(\frac{n_{1}-f}{n_{1}-f+n_{2}}\right)^{2}\sigma^{2}\,g^{-1}\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}};\frac{|\mu_{1}-\mu_{2}|^{2}}{2\sigma^{2}}\Bigr{)}^{2}$
$\quad+\left(\frac{n_{2}}{n_{1}-f+n_{2}}\right)^{2}\sigma^{2}{\frac{4\ln(4/\delta)}{n_{2}}}.$
Rearranging terms, using that $f\geq 0$, and dividing by $4\sigma^{2}$ along with (28) yields that with probability $1-\delta$ we have
$\mathrm{KL}(p_{1}\parallel p)$ $=\frac{|\mu_{1}-\mu|^{2}}{2\sigma^{2}}\geq\frac{|\mu_{1}-\mu_{2}|^{2}}{4\sigma^{2}}-\frac{1}{2}\left(\frac{n_{1}-f}{n_{2}}\right)^{2}g^{-1}\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}};\frac{|\mu_{1}-\mu_{2}|^{2}}{2\sigma^{2}}\Bigr{)}^{2}-{\frac{\ln(4/\delta)}{n_{2}}}$
$=\frac{1}{2}\mathrm{KL}(p_{1}\parallel p_{2})-\frac{1}{2}\left(\frac{n_{1}-f}{n_{2}}\right)^{2}g^{-1}\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}};\frac{|\mu_{1}-\mu_{2}|^{2}}{2\sigma^{2}}\Bigr{)}^{2}-{\frac{\ln(4/\delta)}{n_{2}}}.$
Now, using (28) we get
$\mathrm{KL}(p_{1}\parallel p)\geq\frac{1}{2}\mathrm{KL}(p_{1}\parallel p_{2})-\frac{1}{2}\left(\frac{n_{1}-f}{n_{2}}\right)^{2}g^{-1}\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}};\mathrm{KL}(p_{1}\parallel p_{2})\Bigr{)}^{2}-{\frac{\ln(4/\delta)}{n_{2}}}$
Conclusion.
With probability $1-\delta$, we have $(\alpha,\varepsilon)$-distributional unlearning with
$\alpha$ $\geq\frac{1}{2}\mathrm{KL}(p_{1}\parallel p_{2})-\frac{1}{2}\left(\frac{n_{1}-f}{n_{2}}\right)^{2}g^{-1}\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}};\mathrm{KL}(p_{1}\parallel p_{2})\Bigr{)}^{2}-{\frac{\ln(4/\delta)}{n_{2}}},$
$\varepsilon$ $\leq\left(\frac{n_{1}-f}{n_{2}}\right)^{2}g^{-1}\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}};\mathrm{KL}(p_{1}\parallel p_{2})\Bigr{)}^{2}+\frac{2\ln(4/\delta)}{n_{2}}.$
A.5 Simplified Sample Complexity for Selective Removal
In this section, we can simplify the result of Theorem 2 on selective removal by simplifying cumbersome terms. This leads to Corollary 8, which then yields to the sample complexity results in Table 1.
We first prove an upper bound on the inverse CDF of a folded Normal for small quantiles.
Lemma 7.
Let $\mu_{1},\mu_{2}\in\mathbb{R},\sigma>0$, and $x\sim\mathcal{N}(\mu_{1},\sigma^{2})$. Define the random variable $z=|x-\mu_{2}|$, which follows a folded normal distribution whose cumulative distribution function (CDF) is given by
$F(t)\coloneqq\mathbb{P}[z\leq t]=\Phi\Bigl{(}\frac{t-|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}-\Phi\Bigl{(}\frac{-t-|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)},\quad t\geq 0,$ (32)
where $\Phi$ denotes the standard normal CDF. Then, for any $p>0$ such that $F^{-1}(p)\leq|\mu_{1}-\mu_{2}|$, it holds that:
$F^{-1}(p)\leq\frac{\sigma\,p}{\varphi\Bigl{(}\frac{|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}},$
where $\varphi(x)\coloneqq\frac{1}{\sqrt{2\pi}}\exp\Bigl{(}-\frac{x^{2}}{2}\Bigr{)},x\in\mathbb{R}$, is the standard normal density.
Proof.
Since $F$, defined in (32), is continuously differentiable and strictly increasing on $[0,|\mu_{1}-\mu_{2}|]$ (with $F(0)=0$) and by the assumption $F^{-1}(p)\leq|\mu_{1}-\mu_{2}|$, the Mean Value Theorem guarantees that there exists some $\xi\in[0,F^{-1}(p)]$ such that
$F\Bigl{(}F^{-1}(p)\Bigr{)}=F(0)+F^{\prime}(\xi)\Bigl{(}F^{-1}(p)-0\Bigr{)}.$
Since $F(0)=0$ and $F$ is strictly increasing, one may directly write, via the Mean Value Theorem, that there exists $\xi\in[0,F^{-1}(p)]$ with
$F\Bigl{(}F^{-1}(p)\Bigr{)}=F^{\prime}(\xi)\,F^{-1}(p).$
By definition of the inverse CDF, $F\Bigl{(}F^{-1}(p)\Bigr{)}=p$; hence,
$p=F^{\prime}(\xi)\,F^{-1}(p).$
It remains to lower-bound $F^{\prime}(\xi)$ for $\xi\in[0,|\mu_{1}-\mu_{2}|]$. We recall that for $t\geq 0$,
$F(t)=\Phi\Bigl{(}\frac{t-|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}-\Phi\Bigl{(}\frac{-t-|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}.$
Taking the derivative with respect to $t$ yields
$F^{\prime}(t)$ $=\frac{1}{\sigma}\,\varphi\Bigl{(}\frac{t-|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}+\frac{1}{\sigma}\,\varphi\Bigl{(}\frac{-t-|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}$
$=\frac{1}{\sigma}\,\varphi\Bigl{(}\frac{t-|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}+\frac{1}{\sigma}\,\varphi\Bigl{(}\frac{t+|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)},$
where we used that the standard normal density $\phi$ is symmetric.
For $t\in[0,|\mu_{1}-\mu_{2}|]$, observe that $t-|\mu_{1}-\mu_{2}|\leq 0$. Because the standard normal density is symmetric and nonincreasing on $[0,\infty)$, we have
$\varphi\Bigl{(}\frac{t-|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}=\varphi\Bigl{(}\frac{|\mu_{1}-\mu_{2}|-t}{\sigma}\Bigr{)}\geq\varphi\Bigl{(}\frac{|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}.$
Also, the second term $\varphi\Bigl{(}\frac{t+|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}$ is nonnegative. Hence, for all $t\in[0,|\mu_{1}-\mu_{2}|]$ we have
$F^{\prime}(t)\geq\frac{1}{\sigma}\,\varphi\Bigl{(}\frac{|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}.$
In particular, at $t=\xi$ we obtain
$F^{\prime}(\xi)\geq\frac{1}{\sigma}\,\varphi\Bigl{(}\frac{|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}.$
Substituting this lower bound into the equation $p=F^{\prime}(\xi)\,F^{-1}(p)$ yields
$p\geq\frac{F^{-1}(p)}{\sigma}\,\varphi\Bigl{(}\frac{|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}.$
Rearranging the inequality gives the desired upper bound:
$F^{-1}(p)\leq\frac{\sigma\,p}{\varphi\Bigl{(}\frac{|\mu_{1}-\mu_{2}|}{\sigma}\Bigr{)}}.$
This completes the proof. ∎
We are now ready to prove the corollary below. Note that retaining dependences on $\alpha,\varepsilon$ only in this corollary leads to the result of Table 1.
Corollary 8.
Let $p_{1},p_{2}\in\mathcal{P}$ and $\delta\in(0,1)$. Let $f$ samples from $p_{1}$ be removed according to our distance-based scoring rule or at random before MLE. Then in each case, with probability at least $1-\delta$, the resulting estimate satisfies $(\alpha,\varepsilon)$-distributional unlearning if:
  1. 1.
    Random removal: $n_{2}\geq\tfrac{12\ln(4/\delta)}{\min{\left\{\varepsilon,\alpha\right\}}}$ and $\mathrm{KL}(p_{1}\parallel p_{2})\geq 8\alpha$, and
    $f$ $\geq n_{1}-n_{2}\sqrt{\frac{2\mathrm{KL}(p_{1}\parallel p_{2})-\alpha}{12\mathrm{KL}(p_{1}\parallel p_{2})}},$ $f$ $\geq n_{1}-n_{2}\min{\left\{1,\sqrt{\frac{\varepsilon}{6\mathrm{KL}(p_{1}\parallel p_{2})}}\right\}}.$
  2. 2.
    Selective removal: $n_{2}\geq 2\ln(4/\delta)\max{\left\{\tfrac{1}{\varepsilon},\frac{1}{\sqrt{\varepsilon}},\tfrac{1}{\alpha},\sqrt{\mathrm{KL}(p_{1}\parallel p_{2})-4\alpha}\right\}}$, $\mathrm{KL}(p_{1}\parallel p_{2})\geq 4\alpha$, and $f\geq n_{1}\left(\tfrac{3}{2}+\sqrt{\tfrac{\ln(4/\delta)}{2n_{1}}}-\Phi(2\sqrt{2\mathrm{KL}(p_{1}\parallel p_{2})})\right)$, and
    $f$ $\geq n_{1}-\sqrt{n_{1}n_{2}}\left(\frac{\varepsilon}{16\pi}\right)^{1/4}\exp(-\mathrm{KL}(p_{1}\parallel p_{2})),$
    $f\geq n_{1}-\sqrt{n_{1}n_{2}}\left(\frac{\mathrm{KL}(p_{1}\parallel p_{2})-4\alpha}{8\pi}\right)^{1/4}\exp(-\mathrm{KL}(p_{1}\parallel p_{2})).$
Proof.
We treat random and selective removal separately below.
Random Removal.
Consider removing $f$ samples using the random removal mechanism, before maximum likelihood estimation. From Theorem 1, with probability $1-\delta$, we achieve $(\alpha,\varepsilon)$-unlearning with:
[…]
Therefore, assuming that $n_{2}\geq\tfrac{12\ln(4/\delta)}{\min{\left\{\varepsilon,\alpha\right\}}}$ and $\mathrm{KL}(p_{1}\parallel p_{2})\geq 8\alpha$, direct calculations show that it is sufficient to set:
[…]
Selective Removal.
For selective removal, Theorem 2 shows that, with probability $1-\delta$, we achieve $(\alpha,\varepsilon)$-unlearning with:
[…]
We can simplify these bounds using Lemma 7. Indeed, thanks to the latter and using the same notation and a simple change of variable, we have for all $p,\kappa>0$ such that $p\leq F(|\mu_{1}-\mu_{2}|)$
$g^{-1}\Bigl{(}p;\kappa\Bigr{)}\leq\frac{p}{\phi(\sqrt{2\kappa})}=p\sqrt{2\pi}\exp(\kappa).$ (33)
Now, we plug in $p=1-\tfrac{f}{n_{1}}+\sqrt{\tfrac{\ln(4/\delta)}{2n_{1}}}$. Assuming $f\geq n_{1}\left(-\Phi(2\sqrt{2\mathrm{KL}(p_{1}\parallel p_{2})})+\tfrac{3}{2}+\sqrt{\tfrac{\ln(4/\delta)}{2n_{1}}}\right)$, which directly implies that $p\leq F(|\mu_{1}-\mu_{2}|)$ as required, we then obtain
$g^{-1}\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}};\mathrm{KL}(p_{1}\parallel p_{2})\Bigr{)}\leq\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}}\Bigr{)}\sqrt{2\pi}\exp\left(\mathrm{KL}(p_{1}\parallel p_{2})\right).$
Plugging the above back in the first bounds due to Theorem 2, we obtain:
$\alpha$ $\geq\frac{1}{2}\mathrm{KL}(p_{1}\parallel p_{2})-\frac{1}{2}\left(\frac{n_{1}-f}{n_{2}}\right)^{2}\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}}\Bigr{)}^{2}2\pi\exp\left(2\mathrm{KL}(p_{1}\parallel p_{2})\right)-{\frac{\ln(4/\delta)}{n_{2}}},$
$\varepsilon$ $\leq\left(\frac{n_{1}-f}{n_{2}}\right)^{2}\Bigl{(}1-\frac{f}{n_{1}}+\sqrt{\frac{\ln(4/\delta)}{2n_{1}}}\Bigr{)}^{2}2\pi\exp\left(2\mathrm{KL}(p_{1}\parallel p_{2})\right)+\frac{2\ln(4/\delta)}{n_{2}}.$
Therefore, assuming that $n_{2}\geq 2\ln(4/\delta)\max{\left\{\tfrac{1}{\varepsilon},\frac{1}{\sqrt{\varepsilon}},\tfrac{1}{\alpha},\sqrt{\mathrm{KL}(p_{1}\parallel p_{2})-4\alpha}\right\}}$ and $\mathrm{KL}(p_{1}\parallel p_{2})\geq 4\alpha$, and recalling we had assumed $f\geq n_{1}\left(-\Phi(2\sqrt{2\mathrm{KL}(p_{1}\parallel p_{2})})+\tfrac{3}{2}+\sqrt{\tfrac{\ln(4/\delta)}{2n_{1}}}\right)$, direct calculations show that it is sufficient to set:
[…]
Figure 6: Synthetic Gaussians. Comparison of all strategies considered in the real-world datasets, using the low-divergence scenario of synthetic Gaussians (left plot, Fig. 2). Likelihood-ratio surpasses the distance-based strategy for lower deletion budgets. However, for larger budgets, it sacrifices utility for excessive removal. This is likely because it deletes samples representative of $p_{1}$ but close to $p_{2}$ since $p_{1}$ and $p_{2}$ are similar here.
Appendix B Experimental Details
B.1 Heuristic Deletion Strategies
Across all datasets, we evaluate five scoring strategies for ranking samples in the forget distribution $p_{1}$ for removal. These scores approximate statistical dissimilarity from the retained distribution $p_{2}$ and correspond to different operational interpretations of divergence:
  • lr-cos / lr-maha (likelihood-ratio inspired):
    A proxy for the log-likelihood ratio of $x$ under $p_{2}$ versus $p_{1}$:
    $s(x)=d(x,\mu_{2})-d(x,\mu_{1})$
    where $d(\cdot,\mu)$ denotes cosine distance in TF–IDF space (text) or Mahalanobis distance in ResNet feature space (images). Points are scored high when they are far from $p_{2}$ and close to $p_{1}$.
  • cos-mu2 / maha-mu2 (dissimilarity to $p_{2}$):
    Measures the distance from each $x\in p_{1}$ to the empirical mean of $p_{2}$:
    $s(x)=d(x,\mu_{2})$
    This approximates the contribution of each point to the KL divergence $\mathrm{KL}(p_{2}\|\hat{p})$ when $p$ is modeled as a Gaussian.
  • knn-ratio (local density ratio):
    Estimates the ratio of $k$-NN densities:
    $s(x)=\frac{\widehat{p_{1}}(x)}{\widehat{p_{2}}(x)}$
    where $\widehat{p_{i}}(x)=\exp(-\|x-\text{NN}_{k}^{(i)}(x)\|^{2}/\sigma^{2})$ is a local Gaussian kernel density using $k=10$ nearest neighbors. This captures how typical $x$ is under $p_{1}$ versus $p_{2}$.
  • tfidf-norm / l2-norm:
    Uses the $\ell_{2}$ norm of the raw input (TF–IDF or real-valued) as a proxy for informativeness or deviation from the origin:
    $s(x)=\|x\|_{2}$
  • random:
    Samples points uniformly at random from $p_{1}$ as a baseline.
B.2 Synthetic Gaussians
We draw $n_{1}=n_{2}=1{,}000$ samples from $p_{1}=\mathcal{N}(0,1)$ and $p_{2}=\mathcal{N}(\mu_{2},1)$ for $\mu_{2}\in\{0.5,2.5,5.0\}$, with 20 seeds. After computing scores using each strategy, we remove the top-$f$ fraction of $p_{1}$ points, fit a Gaussian $\mathcal{N}(\hat{\mu},1)$ to the retained data, and compute:
$\alpha=\mathrm{KL}(p_{1}\|\hat{p}),\qquad\varepsilon=\mathrm{KL}(p_{2}\|\hat{p})$
These metrics match the forward-KL objectives of removal and preservation. No predictive model is trained; results reflect pure distributional divergence. For completeness, in Figure 6, we plot a comparison of all strategies considered in the real-world datasets, using the low-divergence scenario of synthetic Gaussians (left plot, Fig. 2). Likelihood-ratio surpasses the distance-based strategy for lower deletion budgets. However, for larger budgets, it sacrifices utility for excessive removal. This is likely because it deletes samples representative of $p_{1}$ but close to $p_{2}$ since $p_{1}$ and $p_{2}$ are similar here. This indicates that no removal strategy strictly dominates all others across all divergence (between $p_{1}$ and $p_{2}$) scenarios.
B.3 Jigsaw Toxic Comments
We use the Jigsaw Toxic Comment Classification dataset, with 140K examples filtered to length 5–200 tokens. We define $p_{1}$ as all training comments containing any of the keywords: “f*ck”, “s*it”, “d*mn”, “b*tch”, “a*s”, and $p_{2}$ as the remaining comments. For each of 5 random seeds, we:
  1. 1.
    Stratified-split $p_{1}$, $p_{2}$ into 70/30 train/val.
  2. 2.
    Compute TF–IDF embeddings (40K max features, 1–2 grams, sublinear TF, min_df=5).
  3. 3.
    Score and remove $f$ of $p_{1}$ training points using each heuristic.
  4. 4.
    Downsample $p_{2}$ to 5× the remaining $p_{1}$ size.
  5. 5.
    Train an $\ell_{2}$-regularized logistic regression on the edited data.
  6. 6.
    Evaluate Recall@$p_{1}$ and macro F1@$p_{2}$ on the validation sets.
B.4 SMS Spam Collection
We use the UCI SMS Spam dataset (5574 examples, 13.4% spam). We apply:
  1. 1.
    TF–IDF vectorization (20K features, 1–2 grams, stopword removal).
  2. 2.
    Scoring of spam ($p_{1}$) messages using each heuristic.
  3. 3.
    Removal of top-$f$ fraction of spam for each strategy.
  4. 4.
    Retrain a logistic regression classifier.
  5. 5.
    Evaluate Recall@spam and F1@ham on a held-out 20% test split.
We run 10 seeds and report mean $\pm$ standard error.
B.5 CIFAR-10 Class Removal
We treat the “cat” class as $p_{1}$ and the other 9 classes as $p_{2}$. We use the standard CIFAR-10 split (50K train, 10K test), and proceed as follows:
  1. 1.
    Train a 3-block CNN (Conv–BN–ReLU ×2 + MaxPool, widths 32–64–128, global avg pool + linear head) for 10 epochs on the full training set.
  2. 2.
    Extract features for all training images using the penultimate layer.
  3. 3.
    Compute Mahalanobis distance scores for each cat image ($p_{1}$) using:
    $s_{\text{maha}}(x)=\sqrt{(x-\mu)^{\top}\Sigma^{-1}(x-\mu)}$
    where $\mu$ and $\Sigma$ are estimated from $p_{2}$.
  4. 4.
    Delete the top-$f$ fraction of cat images under each scoring method.
  5. 5.
    Retrain the same CNN architecture on the edited training set.
  6. 6.
    Evaluate:
    $\text{Accuracy}_{\text{cat}},\qquad\text{Accuracy}_{\text{non-cat}}$
    on the test set. Results are averaged over 30 random seeds.
B.6 Computing Environment
All experiments were run on a HPE DL380 Gen10 equipped with two Intel(R) Xeon(R) Platinum 8358P CPUs running at 2.60GHz, 128 GB of RAM, a 740 GB SSD, and two NVIDIA A10 GPUs. Training for vision experiments was implemented in PyTorch, while text-based experiments used Scikit-learn. All experiments were conducted using a single GPU.